### v2

# Rajiv Raman

Abstract
We study the existence and construction of sparse supports for hypergraphs derived from subgraphs of a graph $G$. For a hypergraph $(X,\mathcal{H})$, a support $Q$ is a graph on $X$ s.t. $Q[H]$, the graph induced on vertices in $H$ is connected for every $H\in\mathcal{H}$.
We consider primal, dual, and intersection hypergraphs defined by subgraphs of a graph $G$ that are non-piercing, (i.e., each subgraph is connected, their pairwise differences remain connected).
If $G$ is outerplanar, we show that the primal, dual and intersection hypergraphs admit supports that are outerplanar. For a bounded treewidth graph $G$, we show that if the subgraphs are non-piercing, then there exist supports for the primal and dual hypergraphs of treewidth $O(2^{\textsc{tw}(G)})$ and $O(2^{4\textsc{tw}(G)})$ respectively, and a support of treewidth $2^{O(2^{\textsc{tw}(G)})}$ for the intersection hypergraph. We also show that for the primal and dual hypergraphs, the exponential blow-up of treewidth is sometimes essential.
All our results are algorithmic and yield polynomial time algorithms (when the treewidth is bounded). The existence and construction of sparse supports is a crucial step in the design and analysis of PTASs and/or sub-exponential time algorithms for several packing and covering problems.
1 Introduction
A hypergraph is defined by a set $X$ of vertices and a collection $\mathcal{H}$ of subsets of $X$, called hyperedges. A support for $(X,\mathcal{H})$ is a graph $Q=(X,F)$ s.t. $Q[H]$ is connected for each $H\in\mathcal{H}$, where $Q[H]$ is the subgraph of $Q$ induced by the vertices in $H$.
Voloshina and Feinberg introduced the notion of a support as a notion of planarity of hypergraphs [voloshina1984planarity]. The notion is now well studied in hypergraph visualization [bereg2015colored, bereg2011red, brandes2010blocks, brandes2012path, buchin2011planar, havet2022overlaying, hurtado2018colored] and network design [anceaume2006semantic, baldoni2007tera, baldoni2007efficient, chand2005semantic, hosoda2012approximability, korach2003clustering, onus2011minimum]. The notion of supports and similar notions have also found applicability for geometric problems [Mustafa17, BasuRoy2018, Cohen-AddadM15, krohn2014guarding, mustafa2010improved, RR18].
Every hypergraph trivially admits a support by putting a complete graph on the vertices. However, we are interested in the existence and construction of sparse supports. Johnson and Pollak [Johnson1987Hypergraph] showed that it is NP-complete to decide if a hypergraph admits a support that is a planar graph.
We consider hypergraphs defined by subgraphs of a graph. Let $G=(V,E)$ be a given graph and let $\mathbf{b}(V)$ and $\mathbf{r}(V)$ be a partition of $V(G)$ into a set of terminals and a set of non-terminals, respectively. A family $\mathcal{H}$ of subgraphs of $G$ defines a hypergraph $(\mathbf{b}(V),\mathcal{H})$, whose vertex set is $\mathbf{b}(V)$ and each $H\in\mathcal{H}$ defines a hyperedge consisting of the vertices in $\mathbf{b}(V)\cap V(H)$. We call this hypergraph the primal hypergraph. Similarly, we define the dual hypergraph whose elements are the subgraphs in $\mathcal{H}$ and each $v\in\mathbf{b}(V)$ defines a hyperedge consisting of all $H\in\mathcal{H}$ containing $v$. We also consider a generalization of the primal and dual hypergraphs namely, the intersection hypergraphs. Let $\mathcal{H}$ and $\mathcal{K}$ be two families of subgraphs of $G$. The intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ is the hypergraph (,$(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}}),\mbox{ where }\mathcal{H}_{K}=\{H\in\mathcal{H}:V(H)\cap V(K)\neq\emptyset\}$
The intersection hypergraphs are a common generalization of the primal and dual hypergraphs. Indeed, setting $\mathcal{H}=\mathbf{b}(V)$, we obtain the primal hypergraphs, and setting $\mathcal{K}=V$, we obtain the dual hypergraphs. Thus, any results that hold for intersection hypergraphs also hold for primal and dual hypergraphs.
A support for a primal (resp. dual/intersection) hypergraph is called a primal (resp. dual/intersection) support. We design algorithms to construct primal, dual and intersection supports for graphs of bounded treewidth. We obtain similar results for a special case of graphs of treewidth 2, namely, the outerplanar graphs.
In the following, the class $\mathcal{G}$ consists either of the family of bounded treewidth graphs, or outerplanar graphs.
{problem}
In order to construct an intersection support in the case of bounded treewidth graphs, we require the construction of primal and dual supports. In the dual setting, note that we can assume without loss of generality that $\mathbf{b}(V)=V$.
As we show below, for the problem defined above, the answer is negative even if $\mathcal{G}$ is the family of trees. However, if we restrict the family of subgraphs to be non-piercing, then the answer to the question above is affirmative.
Definition 1 (Non-piercing).
A family $\mathcal{H}$ of connected subgraphs of a graph $G$ is called non-piercing if the induced subgraph of $G$ on the vertices $V(H)\setminus V(H^{\prime})$ is connected $\forall\;H,H^{\prime}\in\mathcal{H}$.
The notion of non-piercing is motivated by the results of Raman and Ray [RR18], where the authors gave algorithms to construct planar primal, dual and intersection supports for hypergraphs defined by non-piercing regions [^2] A collection $\mathcal{R}$ of connected regions in general position in the plane is called non-piercing if, for each $R,R^{\prime}\in\mathcal{R}$, the region $R\setminus R^{\prime}$ is a connected region in the plane. in the plane. Recently, Raman and Singh [raman2025supportsgraphsboundedgenus] generalized the results of [RR18] to show that if subgraphs of a graph embedded on an oriented surface are cross-free, then there is a support also of the same genus as that of the host graph. Their results imply a unified analysis for several packing and covering problems defined on the hypergraphs. In particular, the existence of a sparse support in many cases implies a PTAS for these problems via a framework of local-search algorithms. See [DBLP:conf/walcom/AschnerKMY13, ChanH12, mustafa2010improved] for a description of the local-search framework. We give two candidate examples where our techniques apply.
Consider the following problem: A collection of cliques in a graph is clearly a collection of non-piercing subgraphs. Given a graph $G$ and a parameter $k\in\mathbb{N}$, find a smallest subset of vertices that hits all cliques of size at least $k$, or a more general version of this problem - Given a set $\mathcal{H}$ of cliques and a set $\mathcal{K}$ of cliques, find a smallest subset of cliques in $\mathcal{H}$ that hit all the cliques in $\mathcal{K}$. As a second example, let $G$ be a graph and let $\mathcal{H}$ be a collection of non-crossing paths in $G$ where we say that two paths $P_{1}$ and $P_{2}$ are non-crossing if once the paths meet at a vertex, they do not separate. Non-crossing paths are, by definition, non-piercing. Suppose we want to select the smallest number of edge-disjoint, or vertex-disjoint paths from $\mathcal{H}$. If $G$ has bounded treewidth, then our results imply the existence of a support of bounded treewidth for the two examples above. Hence, the two problems described above admit a PTAS via a local-search framework.
The non-piercing, or cross-free conditions were used to obtain supports for geometric regions in the plane, or for graphs of bounded genus [raman2025supportsgraphsboundedgenus]. In this paper, since the graph classes are simple, one may wonder if the non-piercing condition is necessary to obtain sparse supports. The following examples show that this is indeed the case. For the primal hypergraph, consider a star $K_{1,n}$ with leaves $v_{1},\ldots,v_{n}$, and $v_{0}$, the central vertex. We color the leaves $\mathbf{b}$ and the central vertex $\mathbf{r}$. The subgraphs $\mathcal{H}$ consist of all pairs of leaves and the central vertex $v_{0}$. Thus, =$\mathcal{H}=\{H_{ij}:H_{ij}=\{v_{i},v_{0},v_{j}\}\}$. It is easy to see that the primal support in this case is $K_{n}$, the complete graph on $n$ vertices. For the dual hypergraph, consider a star $K_{1,\binom{n}{2}}$ with central vertex $v_{0}$. Each leaf $v_{j}$ is labeled by a unique pair $\{x_{j},y_{j}\}$ of $\{1,\ldots,n\}$. There are $n$ subgraphs $H_{1},H_{2},\ldots,H_{n}$, where $H_{i}=\{v_{0}\}\cup\{v_{j}:i\in\{x_{j},y_{j}\}\}$. It is easy to check that the dual support in this case is also $K_{n}$, as there are exactly two subgraphs containing each leaf.
While the notion of supports has found applications in different domains, we still do not have a good tool-set for deciding or constructing supports even in restricted settings, many of which arise in applications. For example, Buchin et al. [buchin2011planar] showed that deciding if a hypergraph admits a 2-outerplanar [^3] A graph $G$ is 2-outerplanar if there is an embedding of $G$ in the plane such that removing the vertices of the outer face results in an outerplanar graph. support is NP-hard, and they left open the question of deciding if a hypergraph admits an outerplanar support. While we do not answer their question, we believe that the tools we develop in constructing an outerplanar support may help in eventually resolving their question.
2 Related Work
The notion of planarity for hypergraphs, unlike that for graphs is not uniquely defined. This notion was first studied by Zykov [zykov], but that was very restrictive. Voloshina and Feinberg [voloshina1984planarity] defined a notion of planarity of hypergraphs that is equivalent to the existence of a support graph that is planar, i.e., a planar support. Johnson and Pollak [Johnson1987Hypergraph] showed that deciding if a hypergraph admits a planar support is NP-hard. Buchin et al. [buchin2011planar] strengthened their result showing that it is NP-hard even to decide if a hypergraph admits a support that is $k$-outerplanar, for $k\geq 2$. However, one can decide in polynomial time if a hypergraph has a support that is a path, a cycle, a cactus [^4] A connected graph is called cactus if each of its edges participates in at most one cycle., or a tree with bounded maximum degree [BOOTH1976335, brandes2010blocks, buchin2011planar, tarjan1984simple].
Besides, other notions of sparsity have also been studied. Du [du1986optimization] proved that it is NP-hard to construct a support with the minimum number of edges.
As stated in the introduction, the existence of a sparse support for an appropriate hypergraph implies a PTAS for several packing and covering problems via a local-search framework. See [BasuRoy2018, ChanH12, mustafa2010improved] for various packing and covering problems in the planar setting including the Independent Set, Dominating Set, Set Cover problems. Raman and Ray [RR18] showed the existence of a planar support for the intersection hypergraph defined by non-piercing regions in the plane. This result was extended by Raman and Singh [raman2025supportsgraphsboundedgenus], who showed that hypergraphs defined by cross-free subgraphs of a graph of genus $g$ also admit a supports of genus at most $g$. Using this, the authors showed that if the hypergraph is defined by a restrictive collection of non-piercing regions on a surface of genus $g$, then there is a support of genus $g$. These results imply a PTAS for all the problems mentioned above.
3 Preliminaries
Let $G=(V,E)$ be a graph and let $\mathcal{H}$ be a collection of connected subgraphs of $G$. For $v\in V$, let $\mathcal{H}_{v}=\{H\in\mathcal{H}:v\in V(H)\}$. Similarly, for an edge $e\in E(G)$, we write $\mathcal{H}_{e}=\{H\in\mathcal{H}:e\in E(H)\}$. We colloquially use $H$ to denote the vertices $V(H)$ and for two subgraphs $X,Y$ of a graph $G$, we write $X\cap Y$ to mean $V(X)\cap V(Y)$, $X\setminus Y$ to mean $V(X)\setminus V(Y)$, and $X\subseteq Y$ to mean $V(X)\subseteq V(Y)$.
Let $\mathbf{b}(V)$ and $\mathbf{r}(V)$ be a partition of $V(G)$ into a set of terminals and a set of non-terminals, respectively. Formally, let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be a 2-coloring of $V(G)$ (not necessarily proper). We call $\mathbf{r}(V)=c^{-1}(\mathbf{r})$ the set of red vertices and $\mathbf{b}(V)=c^{-1}(\mathbf{b})$ the set of blue vertices. We use $\mathbf{b}(H)$ to denote $V(H)\cap\mathbf{b}(V)$, and $\mathbf{r}(H)$ to denote $V(H)\cap\mathbf{r}(V)$. Since two subgraphs are said to intersect if they share a vertex, a hypergraph defined by subgraphs of a graph $G$ remains the same if the subgraphs are replaced by the corresponding induced subgraphs. Moreover, if $\mathcal{H}$ is a collection of non-piercing subgraphs, then it remains non-piercing if the subgraphs are replaced by the corresponding induced subgraphs. Therefore, we assume throughout that the term subgraph refers to an induced subgraph of $G$.
Definition 2 (Outerplanar graph).
A graph $G$ is called outerplanar if there is an embedding of $G$ in the plane such that all its vertices lie on the exterior face.
Next, we define the notion of tree decomposition. Throughout the paper, we use the term node to refer to the elements of $V(T)$ for a tree $T$, and we use vertices to refer to the elements of $V(G)$ for a graph $G$. We use letters $x,y,z$ for the nodes of $T$, and $u,v,w$ for the vertices of $G$.
Definition 3 (Tree decomposition).
Given a graph $G=(V,E)$, a tree decomposition of $G$ is a pair $(T,\mathcal{B})$, where $T$ is a tree and $\mathcal{B}$ is a collection of bags - subgraphs of $G$ indexed by the nodes of $T$, that satisfies the following properties:
  1. 1.
    For each $v\in V(G)$, the set of bags of $T$ containing $v$ induces a sub-tree of $T$.
  2. 2.
    For every edge $\{u,v\}$ in $G$, there is a bag $B\in\mathcal{B}$ such that $u,v\in B$.
Definition 4 (Treewidth).
The width of a tree decomposition $(T,\mathcal{B})$ is defined to be $\max_{x\in V(T)}|B_{x}|-1$. The treewidth of a graph $G$ is the minimum width over all the tree decompositions of $G$, and is denoted $\textsc{tw}(G)$.
We use two well-known properties of tree decompositions. First, that we can assume that $T$ is a rooted binary tree. That is, we can modify a given tree decomposition to one where $T$ is a rooted binary tree without increasing the treewidth. If $T$ is rooted, then we use $T_{x}$ to denote the subtree rooted at node $x$. The second is that for any edge $e=\{x,y\}$ of $T$, the set $B_{x}\cap B_{y}$ is a separator in $G$. That is, the induced subgraph $G^{\prime}=G\setminus\{B_{x}\cap B_{y}\}$ of $G$ is disconnected. The set $B_{x}\cap B_{y}$ is called the adhesion set corresponding to edge $\{x,y\}$ of $T$. We refer the reader to Chapter 12 in [diestel2005graph] for more details on the properties of treewidth and tree decompositions.
For a graph $G$ and collections $\mathcal{H}$ and $\mathcal{K}$ of subgraphs of $G$, we use the pair $(G,\mathcal{H})$ to refer to the primal or dual hypergraph, and triple $(G,\mathcal{H},\mathcal{K})$ for the intersection hypergraph. We call the tuple $(G,\mathcal{H})$ a graph system and the tuple $(G,\mathcal{H},\mathcal{K})$ an intersection system. If $G\in\mathcal{G}$, and $\mathcal{H,K}$ satisfy a property $\Pi$, then we say that $(G,\mathcal{H})$ is a $\mathcal{G}$ $\Pi$ graph system, or that $(G,\mathcal{H},\mathcal{K})$ is a $\mathcal{G}$ $\Pi$ intersection system. For example, if $G$ has bounded treewidth, and $\mathcal{H}$ and $\mathcal{K}$ are connected subgraphs of $G$, then $(G,\mathcal{H},\mathcal{K})$ is a bounded treewidth connected intersection system. If $G$ is outerplanar and $\mathcal{H}$ is non-piercing, then $(G,\mathcal{H})$ is an outerplanar non-piercing system.
Note that adding edges to $G$ leaves the subgraphs in $\mathcal{H}$ or $\mathcal{K}$ non-piercing. Thus, we can assume in the outerplanar setting that $G$ is a maximal outerplanar graph, and in the setting where $G$ has bounded treewidth, we can assume that each bag in a tree decomposition of the graph induces a complete graph, as this does not increase the treewidth (See Section 6 for more details).
Let $\mathcal{H}$ be a collection of subgraphs of $G$. For two subgraphs $H,H^{\prime}\in\mathcal{H}$ if $H\subseteq H^{\prime}$, we say that $H$ is contained in $H^{\prime}$. If there are no such subgraphs $H,H^{\prime}$, we say that $\mathcal{H}$ is containment-free. The containments in $\mathcal{H}$ yields a natural partial order $\prec_{N}$ on $\mathcal{H}$, where $H\prec_{N}H^{\prime}\Leftrightarrow V(H)\subseteq V(H^{\prime})$. Let $\mathcal{H}^{*}\subseteq\mathcal{H}$ be the containment-free subfamily of $\mathcal{H}$ consisting of the maximal elements of this partial order. For an intersection support (and hence for a dual support), the following result shows that it is sufficient to assume that $\mathcal{H}$ is containment-free.
Lemma 5.
Let $(G,\mathcal{H,K})$ be a graph system and $\mathcal{H}^{*}\subseteq\mathcal{H}$ be the containment-free subfamily of $\mathcal{H}$ consisting of all the maximal elements of $(\mathcal{H},\prec_{N})$. Then,
  1. 1.
    If $(G,\mathcal{H^{*},K})$ has an outerplanar support, then $(G,\mathcal{H,K})$ also has an outerplanar support.
  2. 2.
    If $(G,\mathcal{H^{*},K})$ has a support of treewidth $t$, then $(G,\mathcal{H,K})$ also has a support of treewidth $t$.
Proof.
By definition, any subgraph $H\in\mathcal{H}\setminus\mathcal{H}^{*}$ has a successor $H^{\prime}\in\mathcal{H}^{*}$ in the poset $(\mathcal{H},\preceq_{N})$ i.e., $H\subseteq H^{\prime}$ for some $H^{\prime}\in\mathcal{H}^{*}$. Let $\tilde{Q}^{\prime}$ be a support for $(G,\mathcal{H}^{*},\mathcal{K})$, and $\tilde{Q}$ be a graph obtained from $\tilde{Q}^{\prime}$ by adding a vertex for each $H\in\mathcal{H}\setminus\mathcal{H}^{*}$ and making it adjacent to one of its successors $H^{\prime}\in\mathcal{H}^{*}$ (breaking ties arbitrarily). Since $\tilde{Q}$ is obtained form $\tilde{Q}^{\prime}$ by adding only degree one vertices, $\textsc{tw}(\tilde{Q})=\textsc{tw}(\tilde{Q}^{\prime})$. Moreover, $\tilde{Q}$ remains outerplanar if $\tilde{Q}^{\prime}$ is outerplanar.
We now show that $\tilde{Q}$ is an intersection support for $(G,\mathcal{H,K})$. Let $K\in\mathcal{K}$ be arbitrary. We claim that $\tilde{Q}[\mathcal{H}_{K}]$ induces a connected subgraph of $\tilde{Q}$. Since $\tilde{Q}^{\prime}$ is an intersection support for $(G,\mathcal{H}^{*},\mathcal{K})$, the induced subgraph $\tilde{Q}^{\prime}[\mathcal{H}^{*}_{K}]$ of $\tilde{Q}^{\prime}$ is connected. By definition, $\mathcal{H}^{*}\subseteq\mathcal{H}$ consists of precisely the maximal elements of $(\mathcal{H},\preceq_{N})$. Therefore, for any $H\in\mathcal{H}_{K}\setminus\mathcal{H}^{*}_{K}$, $\exists H^{\prime}\in\mathcal{H}^{*}$ such that $H\subseteq H^{\prime}$ and $\{H,H^{\prime}\}$ is an edge in $\tilde{Q}$. But $H^{\prime}$ intersects $K$ since $H$ interacts $K$. It follows that $H^{\prime}\in\mathcal{H}^{*}_{K}$ and hence, the induced subgraph $\tilde{Q}[\mathcal{H}_{K}]$ of $\tilde{Q}$ is connected. This completes the proof.
4 Contributions
We give polynomial time algorithms to construct an outerplanar intersection support for an outerplanar non-piercing system $(G,\mathcal{H},\mathcal{K})$. For a non-piercing graph system of treewidth $t$, we show that the supports for the primal and dual hypergraphs have treewidth $O(2^{t})$ and $O(2^{4t})$ respectively, and the intersection support has treewidth $2^{O(2^{t})}$. For the primal and dual settings, we show that the exponential blow-up in the treewidth of the support is sometimes necessary.
Theorem 6.
If $(G,\mathcal{H,K})$ is an outerplanar non-piercing system, then there is an intersection support for $(G,\mathcal{H,K})$ which is an outerplanar graph.
This also implies the existence of primal and dual supports for an outerplanar non-piercing graph system.
Theorem 7.
Let $(G,\mathcal{H})$ be a non-piercing graph system of treewidth $t$ with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$. There is a primal support $Q$ of treewidth at most $2^{t+2}+t$.
Theorem 8.
Let $(G,\mathcal{H})$ be a non-piercing graph system of treewidth $t$. There is a dual support $Q^{*}$ of treewidth at most $2^{4(t+1)}$.
Theorem 9.
Let $(G,\mathcal{H,K})$ be a non-piercing intersection system of treewidth $t$. Then, there is an intersection support $\tilde{Q}$ of treewidth at most $2^{2^{t+4}+4(t+1)}$.
The rest of the paper is organized as follows. In Section 5, we present our results on outerplanar graphs. In Section 6, we present our results on bounded treewidth graphs. We conclude in Section 8.
5 Outerplanar Support
In this section, we give a polynomial time algorithm that computes an intersection support for an outerplanar non-piercing intersection system.
Recall from Section 3 that we can assume $G$ is maximal outerplanar. Let $C$ be the outer face in an outerplanar embedding of $G$. Let $\mathcal{H},\mathcal{K}$ be two families of connected subgraphs of $G$. Observe that two subgraphs of a graph intersect if and only if they share a vertex. Hence, the hypergraph on $\mathcal{H},\mathcal{K}$ induced on $C$ is identical to that induced on $G$. Consequently, it is sufficient to construct an outerplanar support for the cycle systems that are strong $axax$-free, a notion we define below.
5.1 Cycle $axax$-free systems
For a cycle $C$ embedded in the plane, and collections $\mathcal{H}$ and $\mathcal{K}$ of subgraphs (not necessarily connected) of $C$, we call $(C,\mathcal{H})$ a cycle system and $(C,\mathcal{H},\mathcal{K})$ a cycle intersection system. We start with the notions of cycle systems that (strong) $axax$-free and $abab$-free.
Definition 10 ($axax$-free).
Let $(C,\mathcal{H})$ be a cycle system. $H,H^{\prime}\in\mathcal{H}$ are an $axax$-pair if there are four distinct vertices $a_{1},x_{1},a_{2},x_{2}$ in cyclic order on $C$ such that $a_{1},a_{2}\in H\setminus H^{\prime}$ and $x_{1},x_{2}\in H^{\prime}$. $(C,\mathcal{H})$ is $axax$-free if there are no $axax$-pairs in $\mathcal{H}$.
Definition 11 ($abab$-free).
Let $(C,\mathcal{H})$ be a cycle system. $H,H^{\prime}\in\mathcal{H}$ are an $abab$-pair if there are four distinct vertices $a_{1},b_{1},a_{2},b_{2}$ in cyclic order on $C$ such that $a_{1},a_{2}\in H\setminus H^{\prime}$ and $b_{1},b_{2}\in H^{\prime}\setminus H$. $(C,\mathcal{H})$ is $abab$-free if there are no $abab$-pairs in $\mathcal{H}$.
Clearly, if $(C,\mathcal{H})$ is $axax$-free, then it is $abab$-free. We need one more definition, that of strong $axax$-free cycle systems.
Definition 12 ((Strong) $axax$-free).
  1. 1.
    For two families $\mathcal{H},\mathcal{K}$ of subgraphs of $C$, the intersection system $(C,\mathcal{H},\mathcal{K})$ satisfies the $axax$-free property if both $(C,\mathcal{H})$ and $(C,\mathcal{K})$ are $axax$-free.
  2. 2.
    $(C,\mathcal{H},\mathcal{K})$ is said to satisfy the intersection property if for any $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with four vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order on $C$ such that $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$, then $H\cap K\neq\emptyset$.
  3. 3.
    If $(C,\mathcal{H},\mathcal{K})$ satisfies the $axax$-free property and the intersection property, then it is said to satisfy the strong $axax$-free property.
Lemma 13.
Let $(G,\mathcal{H},\mathcal{K})$ be an embedded outerplanar non-piercing system with $C$ denoting the outer cycle of $G$. Then, $(C,\mathcal{H},\mathcal{K})$ satisfies the strong $axax$-free property.
Proof.
Let $H,H^{\prime}\in\mathcal{H}$. Since $\mathcal{H}$ are non-piercing subgraphs of $G$, $H\setminus H^{\prime}$ and $H^{\prime}\setminus H$ are both connected. Suppose there is a cyclic sequence $a_{1},x_{1},a_{2},x_{2}$ of vertices on $C$ such that $a_{1},a_{2}\in H\setminus H^{\prime}$, and $x_{1},x_{2}\in H^{\prime}$. Since $H\setminus H^{\prime}$ is connected, there is a path $P$ between $a_{1}$ and $a_{2}$ lying in $H\setminus H^{\prime}$. Then, $x_{1}$ and $x_{2}$ lie in separate components of $G\setminus P$, contradicting the assumption that $H^{\prime}\setminus H$ is connected. An identical argument shows that $(C,\mathcal{K})$ is $axax$-free. See Figure 1.
To see that $(C,\mathcal{H},\mathcal{K})$ satisfies the intersection property, consider $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order s.t. $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$. Since $H$ is connected in $G$, there is a path $P$ between $h_{1}$ and $h_{2}$, all of whose vertices lie in $H$. If $H\cap K=\emptyset$, then $P$ lies in $H\setminus K$. Since $h_{1}$ and $h_{2}$ are non-consecutive on $C$, $G\setminus P$ gets separated into two components with $k_{1}$ and $k_{2}$ in distinct components. This implies there is no path between $k_{1}$ and $k_{2}$ in $G$ that lies entirely in $K$, a contradiction since $K$ is connected.
Figure 1: Non-piercing implies $axax$-free. The figure is from [raman2025supportsgraphsboundedgenus].
Our next theorem shows that if $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free, then it admits an outerplanar intersection support. We will use this to prove the existence of an outerplanar support for the non-piercing intersection system $(G,\mathcal{H},\mathcal{K})$.
Theorem 14.
Let $(C,\mathcal{H},\mathcal{K})$ be a strong $axax$-free outerplanar system. Then, there is an outerplanar support $Q$ for $(C,\mathcal{H},\mathcal{K})$.
Now, Theorem 6 is a direct implication of Lemma 13 and Theorem 14.
See 6
Proof.
Let $C$ be the outer cycle in an outerplanar embedding of $G$. By Lemma 13, $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. By Theorem 14, there is an outerplanar support $\tilde{Q}$ for $(C,\mathcal{H},\mathcal{K})$. $\tilde{Q}$ is also a support for $(G,\mathcal{H},\mathcal{K})$ since the underlying intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ remains the same on $G$ and on $C$.
For Theorem 6 to go through, it remains to prove Theorem 14 which we do in Section 5.2. Below, we show that we can simplify an $axax$-free cycle system so that each vertex of $C$ is contained in at least one subgraph in $\mathcal{H}$ and one subgraph in $\mathcal{K}$.
Definition 15 (Reduced cycle system).
A cycle system $(C,\mathcal{H},\mathcal{K})$ is called reduced if for each vertex $v$ of $C$, there is an $H\in\mathcal{H}$ and a $K\in\mathcal{K}$ such that $v\in V(H)\cap V(K)$.
Consider a cycle system $(C,\mathcal{H,K})$. Let $C^{\prime}$ be the induced cycle obtained from $C$ by removing each vertex $v$ of $C$ and making its neighbours adjacent, whenever $v\notin V(H)$ for all $H\in\mathcal{H}$ or $v\notin V(K)$ for all $K\in\mathcal{K}$. Let =$\mathcal{H}^{\prime}=\{H\cap C^{\prime}:H\in\mathcal{H}\}$ and $\mathcal{K}^{\prime}=\{K\cap C^{\prime}:K\in\mathcal{K}\}$. It follows that $(C^{\prime},\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ is a reduced cycle system. The following proposition shows that it is sufficient to construct a support when $(C,\mathcal{H,K})$ is a reduced system.
Proposition 16.
If $(C,\mathcal{H,K})$ is a strong $axax$-free system, then the reduced system $(C^{\prime},\mathcal{H^{\prime},K^{\prime}})$ is also a strong $axax$-free system. Further, any intersection support for $(C^{\prime},\mathcal{H^{\prime},K^{\prime}})$ is also an intersection support for $(C,\mathcal{H,K})$.
Proof.
If $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free, removing an $\mathcal{H}$-vertex or a $\mathcal{K}$-vertex leaves the reduced system $(C^{\prime},\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ is strong $axax$-free and the resulting intersection hypergraph remains the same since we did not remove any vertex $v\in H\cap K$ for any $H\in\mathcal{H}$ and $K\in\mathcal{K}$. Hence, a support for the reduced system is also a support for the original system.
As a consequence of Proposition 16, we assume throughout that $(C,\mathcal{H},\mathcal{K})$ is a reduced system.
5.2 Construction of Outerplanar Supports
In this section, we prove Theorem 14. We start with some basic terminology required for the proof. Let $C=\{0,\ldots,n-1\}$ be a cycle on $n$ vertices oriented clockwise, and let $\mathcal{R}$ be a collection of arcs on $C$ whose both ends are defined by the vertices of $C$ such that no two arcs in $\mathcal{R}$ contain the same set of vertices.
For an $R\in\mathcal{R}$, if $R=\mathrm{arc}[i,j]$, i.e., $R$ consists of a consecutive sequence of vertices, i.e., $R=[i,i+1,\ldots,j]$ where the indices are numbered $\pmod{n}$, we say that $R$ is a run on $C$. We also use $\mathrm{arc}(i,j)$ to denote the run $[i+1,\ldots,j-1]$. Let $s(R)=i$ and $t(R)=j$. Consider a pair of arcs $R,R^{\prime}\in\mathcal{R}$ such that $R\subseteq R^{\prime}$. In a traversal of $C$ starting at $s(R^{\prime})$, if we have $s(R^{\prime})<s(R)<t(R)<t(R^{\prime})$, then we say that $R$ is strictly contained in $R^{\prime}$. If $s(R)=s(R^{\prime})$, or $t(R)=t(R^{\prime})$, then we say that $R$ is weakly contained in $R^{\prime}$. If there is no pair of arcs $R,R^{\prime}\in\mathcal{R}$ such that $R$ is strictly contained in $R^{\prime}$, then we say that $\mathcal{R}$ is strict-containment free. If there is no pair of arcs $R,R^{\prime}\in\mathcal{R}$ that are weakly contained in each other or strictly contained in each other, then $\mathcal{R}$ is containment-free i.e., $R\setminus R^{\prime}\neq\emptyset$ for all $R,R^{\prime}\in\mathcal{R}$. See Figure 2.
((a)) $R_{2}$ is weakly contained in $R_{1}$, and $R_{4}$ is strictly contained in $R_{3}$.
((b)) Strict-containment free
Figure 2: Weak containment, strict containment and strict-containment free. Induced subgraphs of the cycle are shown by corresponding arcs inside the cycle.
Strict containment defines a natural partial order on $\mathcal{R}$. Let $\mathcal{R}^{*}\subseteq\mathcal{R}$ denote the maximal elements of this strict-containment order. Then, $\mathcal{R^{*}}$ is a maximal strict-containment free subset of $\mathcal{R}$. Note that $\mathcal{R}^{*}$ can contain subgraphs that are weakly contained in each other.
An immediate consequence of a cycle system $(C,\mathcal{H},\mathcal{K})$ being $axax$-free that will be useful later is the following proposition which is a slight variation of Lemma 5. The proof is along the same lines as of Lemma 5.
Proposition 17.
Let $(C,\mathcal{H,K})$ be a cycle $axax$-free system where subgraphs in $\mathcal{H}$ are single runs, and let $\mathcal{H}^{*}$ be the maximal elements in the strict-containment order on $\mathcal{H}$. If $(C,\mathcal{H^{*},K})$ admits an outerplanar support, then so does $(C,\mathcal{H,K})$.
As a consequence of Lemma 5, or Proposition 17, we can assume that $\mathcal{H}$ is either containment-free, or it is strict-containment free (as will be useful in our proofs).
Let $\mathcal{R}$ be a collection of arcs on a cycle $C$ that are strict-containment free. Starting at an arbitrary vertex on $C$, in a clockwise traversal, we construct a cycle $C(\mathcal{R})$ on $\mathcal{R}$ as follows: When we visit a vertex $v$ of $C$, for each $R\in\mathcal{R}$ s.t. $s(R)=v$, we put a vertex in $C(\mathcal{R})$. The vertices added to $C(\mathcal{R})$ w.r.t. $v$ are ordered in increasing order of $t(R)$, with ties broken arbitrarily. We call the cycle $C(\mathcal{R})$ thus constructed, the cycle in lex. cyclic order, and we say that $\mathcal{R}$ is ordered in lex. cyclic order.
In a cycle system $(C,\mathcal{H},\mathcal{K})$, each subgraph $X\in\mathcal{H}\cup\mathcal{K}$ induces a sequence of runs on $C$. Let $n_{X}$ denote the number of runs of $X$ in $C$. Let $r_{0}(X)=[s_{0},\ldots,t_{0}],r_{1}(X)=[s_{1},\ldots,t_{1}],\ldots,r_{k-1}=[s_{k-1},\ldots,t_{k-1}]$ denote the $k=n_{X}$ runs of $X$ on $C$. For $i=0,\ldots,k-1$, let $d_{i}$ denote the chord $\{t_{i},s_{i+1}\}$. Let $\ell(d_{i})=|\mathrm{arc}[t_{i},s_{i+1}]|$, and let $\ell(X)=\arg\min_{d_{i}}\ell(d_{i})$, with ties broken arbitrarily. If $n_{X}=1$, then $d_{X}$ is undefined.
The proof of Theorem 14 proceeds in three easy steps. We start with the case when all subgraphs in $\mathcal{H}\cup\mathcal{K}$ induce single runs on $C$. We next use this result to obtain an outerplanar support for the case when only the subgraphs in $\mathcal{H}$ are required to induce single runs on $C$. Finally, we use the outerplanar supports guaranteed by the two special cases to obtain an outerplanar support for the general case.
Let each subgraph in $\mathcal{H}\cup\mathcal{K}$ induce a single run on $C$. For the requirement in the general settings, we assume by Proposition 17 that $\mathcal{H}$ is strict-containment free. We claim that $C(\mathcal{H})$, the cycle on $\mathcal{H}$ in lex. cyclic order is the desired support for $(C,\mathcal{H},\mathcal{K})$.
Lemma 18.
Let $(C,\mathcal{H,K})$ be a cycle system such that $\mathcal{H}$ is strict-containment free and each $X\in\mathcal{H\cup K}$ induce a single run on $C$. Then, the cycle $C(\mathcal{H})$ is an outerplanar support for $(C,\mathcal{H,K})$.
Proof.
$C(\mathcal{H})$ is clearly an outerplanar graph. Consider a subgraph $K\in\mathcal{K}$. Since $K$ induces a single run on $C$ and the subgraphs in $\mathcal{H}$ are strict-containment free, it follows that the subgraphs in $\mathcal{H}_{K}$ appear consecutively in lex. cyclic order on $C(\mathcal{H})$ and thus induce a connected subgraph of $C(\mathcal{H})$.
Now assume that subgraphs in $\mathcal{H}$ induce single runs on $C$, and those in $\mathcal{K}$ can have multiple runs. We claim that there is an outerplanar support with outer cycle $C(\mathcal{H})$. But, before we do that, we start with the following simple consequence of $(C,\mathcal{H},\mathcal{K})$ being $axax$-free.
Proposition 19.
Let $(C,\mathcal{H})$ be $axax$-free. For an $H\in\mathcal{H}$, let $u,v$ be two non-consecutive vertices on $C$ with $u,v\in H$. Then, for any $H^{\prime}\in\mathcal{H}$ s.t. $H^{\prime}\cap\mathrm{arc}(u,v)\neq\emptyset\neq H^{\prime}\cap\mathrm{arc}(v,u)$, $H^{\prime}$ must contain either $u$ or $v$. Moreover, if $H\cap\mathrm{arc}(u,v)=\emptyset$, then $H^{\prime}\cap\mathrm{arc}[v,u]\subseteq H\cap\mathrm{arc}[v,u]$.
Proof.
Let $x$ and $y$ be any two vertices of $H^{\prime}$ in $\mathrm{arc}(u,v)$ and $\mathrm{arc}(v,u)$ respectively, and let $u,v\in H\setminus H^{\prime}$. The vertices $u,x,v,y$ form an $axax$-pattern on $C$; a contradiction since $(C,\mathcal{H})$ is $axax$-free.
The proof for the moreover part is similar. Indeed, let $h^{\prime}\in\mathrm{arc}[v,u]$ s.t. $h^{\prime}\in H^{\prime}\setminus H$. Also, note that $x\in H^{\prime}\setminus H$. Hence, the subgraphs $H,H^{\prime}$ form an $axax$-pattern as witnessed by the cyclic sequence $h^{\prime},u,x,v$; a contradiction.
Lemma 20.
Let $(C,\mathcal{H,K})$ be any $axax$-free cycle system s.t. each $H\in\mathcal{H}$ induces a single run on $C$. If $\mathcal{H}$ is strict-containment free, then there exists an outerplanar support for $(C,\mathcal{H,K})$ with outer cycle $C(\mathcal{H})$.
Proof.
Let $N=N(C,\mathcal{K})=\sum_{K\in\mathcal{K}}(n_{K}-1)$, where $n_{K}$ is the number of runs of $K$ on $C$. We proceed by induction on $N$. If $N=0$, then $n_{K}=1$ for all $K\in\mathcal{K}$ and we are done by Lemma 18. So, suppose $N\geq 1$ i.e., $n_{K}\geq 2$ for some $K\in\mathcal{K}$.
We assume the lemma holds when $N(C,\mathcal{K}^{\prime})<N$ for any $axax$-free system $(C,\mathcal{H,K^{\prime}})$. Let $(C,\mathcal{H,K})$ be an $axax$-free system with $N(C,\mathcal{K})=N$. Let $K_{0}=\arg\min_{K\in\mathcal{K}}\ell(K)$. Let $d_{K_{0}}=\{u_{0},v_{0}\}$ be the chord of $K_{0}$ realizing this minimum. We use the chord $d_{K_{0}}$ to split $C$ into two cycles.
Let $C_{L}$ denote the cycle $\mathrm{arc}[v_{0},u_{0}]\cup d_{K_{0}}$, and $C_{R}$ denote the cycle $\mathrm{arc}[u_{0},v_{0}]\cup d_{K_{0}}$. Let $\mathcal{K}_{X}=\{K\cap C_{X}:K\in\mathcal{K}\}$, where $X\in\{L,R\}$. Observe that $K_{0}$ appears both in $C_{L}$ and $C_{R}$, where in $C_{R}$, $K_{0}$ spans the vertices $\{u_{0},v_{0}\}$.
Let $H_{1},\ldots,H_{r}$ be the subgraphs in $\mathcal{H}_{u_{0}}$ in lex. cyclic order, i.e., labeled such that $(H_{r}\cap C_{L})\subseteq(H_{r-1}\cap C_{L})\subseteq\ldots\subseteq(H_{1}\cap C_{L})$. Since $\mathcal{H}$ is strict-containment free, it follows that $(H_{1}\cap C_{R})\subseteq(H_{2}\cap C_{R})\subseteq\ldots\subseteq(H_{r}\cap C_{R})$. Similarly, if $H^{\prime}_{1},\ldots,H^{\prime}_{s}$ are the subgraphs in $\mathcal{H}_{v_{0}}$ in lex. cyclic order, then $(H^{\prime}_{1}\cap C_{L})\subseteq(H^{\prime}_{2}\cap C_{L})\subseteq\ldots\subseteq(H^{\prime}_{s}\cap C_{L})$, and $(H^{\prime}_{s}\cap C_{R})\subseteq(H^{\prime}_{s-1}\subseteq C_{R})\ldots\subseteq(H^{\prime}_{1}\cap C_{R})$. Figure 3 shows the subgraphs in $\mathcal{H}_{u_{0}}$ and in $\mathcal{H}_{v_{0}}$.
We let $\mathcal{H}_{L}=\{H\in\mathcal{H}:H\subseteq\mathrm{arc}(v_{0},u_{0})\}\cup\{H_{1}\cap C_{L}\}\cup\{H^{\prime}_{s}\cap C_{L}\}$, and $\mathcal{H}_{R}=\{H\cap C_{R}:H\in\mathcal{H}\}$. See Figure 4.
Figure 3: Ordering $\mathcal{H}_{u_{0}}$ and $\mathcal{H}_{v_{0}}$ subgraphs in $C_{L}$ and $C_{R}$. $(H_{3}\cap C_{L})\subseteq(H_{2}\cap C_{L})\subseteq(H_{1}\cap C_{L})$, and $(H_{1}\cap C_{R})\subseteq(H_{2}\cap C_{R})\subseteq(H_{3}\cap C_{R})$. Analogously, $(H^{\prime}_{1}\cap C_{L})\subseteq(H^{\prime}_{2}\cap C_{L})\subseteq(H^{\prime}_{3}\cap C_{L})$, and $(H^{\prime}_{3}\cap C_{R})\subseteq(H^{\prime}_{2}\cap C_{R})\subseteq(H^{\prime}_{1}\cap C_{R})$.
Figure 4: Construction of $\mathcal{H}_{L}$ and $\mathcal{H}_{R}$. Any $H\in\mathcal{H}\setminus\{H_{1},H^{\prime}_{3}\}$ containing $u_{0}$ or $v_{0}$ such that $(H\cap C_{L})\subseteq(H_{1}\cap C_{L})$ or $(H\cap C_{L})\subseteq(H^{\prime}_{3}\cap C_{L})$ is not included in $\mathcal{H}_{L}$.
Note that since $\mathcal{H}$ is strict-containment free, so are $\mathcal{H}_{L}$ and $\mathcal{H}_{R}$, and that each $H\in\mathcal{H}\setminus\{H_{1},H^{\prime}_{s}\}$ contributes a run to exactly one of $\mathcal{H}_{L}$ or $\mathcal{H}_{R}$, by construction.
By the choice of $K_{0}$, in the cycle system $(C_{R},\mathcal{H}_{R},\mathcal{K}_{R})$, each $H\in\mathcal{H}_{R}$ and each $K\in\mathcal{K}_{R}$ induces a single run on $C_{R}$. Hence, by Lemma 18, there is an outerplanar support $Q_{R}$ for $(C_{R},\mathcal{H}_{R},\mathcal{K}_{R})$ such that $Q_{R}=C(\mathcal{H}_{R})$, the cycle on $\mathcal{H}_{R}$ in lex. cyclic order.
The cycle system $(C_{L},\mathcal{H}_{L},\mathcal{K}_{L})$ is $axax$-free since $(C,\mathcal{H},\mathcal{K})$ is $axax$-free. Further, $N(C_{L},\mathcal{K}_{L})<N$ since $d_{K_{0}}$ joins two disjoint runs of $K_{0}$. Hence, by the inductive hypothesis, $(C_{L},\mathcal{H}_{L},\mathcal{K}_{L})$ admits an outerplanar support $Q_{L}$ with outer cycle $C(\mathcal{H}_{L})$.
Since in $Q_{L}$, the outer cycle is $C(\mathcal{H}_{L})$ and in $Q_{R}$, the outer cycle is $C(\mathcal{H}_{R})$, it follows that $H_{1}$ and $H^{\prime}_{s}$ are consecutive in the outer cycles of $Q_{L}$ and $Q_{R}$. To obtain a support $Q$ for $(C,\mathcal{H,K})$, we identify the copy of $H_{1}$ in $Q_{L}$ and $Q_{R}$, and similarly, we identify the copy of $H^{\prime}_{s}$ in $Q_{L}$ and $Q_{R}$. It follows that $Q$ is outerplanar and $C(\mathcal{H})$ is the outer cycle in the resulting embedding of $Q$.
Next, we show that $Q$ is a support. Consider an arbitrary subgraph $K\in\mathcal{K}$. We consider three possible cases for the runs of $K$.
Suppose the runs of $K$ lie entirely in $C_{R}$. Then, $\mathcal{H}_{K}$ induces a connected subgraph in $Q$ since every subgraph in $\mathcal{H}$ intersecting $C_{R}$ has a run in $\mathcal{H}_{R}$, and that $Q_{R}$ is a support for $(C_{R},\mathcal{H}_{R},\mathcal{K}_{R})$.
Now, suppose the runs of $K$ lie entirely in $C_{L}$. If $\mathcal{H}_{K}\cap(\mathcal{H}_{u_{0}}\cup\mathcal{H}_{v_{0}})=\emptyset$, i.e., if none of the subgraphs in $\mathcal{H}_{K}$ contains $u_{0}$ or $v_{0}$, then for each subgraph $H\in\mathcal{H}_{K}$, we have $H\subseteq\mathrm{arc}(v_{0},u_{0})$ and thus $H\in\mathcal{H}_{L}$ since each $H\in\mathcal{H}$ induces a single run on $C$. Since $Q_{L}$ is a support for $(C_{L},\mathcal{H}_{L},\mathcal{K}_{L})$, it follows that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. Now, suppose $\mathcal{H}_{K}\cap\mathcal{H}_{u_{0}}\neq\emptyset$, or $\mathcal{H}_{K}\cap\mathcal{H}_{v_{0}}\neq\emptyset$. Assume the former, without loss of generality. Since the subgraphs in $\mathcal{H}$ were assumed to be strict-containment free, $K$ intersects a prefix of the sequence of subgraphs in $\mathcal{H}_{u_{0}}$ in lex. cyclic order, i.e., a prefix of the subgraphs $(H_{1},\ldots,H_{r})$, since $(H_{i}\cap C_{L})\supseteq(H_{i+1}\cap C_{L})$ for $i=1,\ldots,r-1$. In particular $H_{1}\in\mathcal{H}_{K}$. Again, as argued, the subgraphs in $\mathcal{H}_{R}$ are strict-containment free, and hence, they appear consecutively in lex. cyclic order on the outer cycle of $Q_{R}$. Since $Q_{R}$ is a support for $(C_{R},\mathcal{H}_{R},\mathcal{K}_{R})$, and $Q_{L}$ is a support for $(C_{L},\mathcal{H}_{L},\mathcal{K}_{L})$, with $H_{1}$ appearing in both $Q_{L}$ and $Q_{R}$, it follows that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. A similar argument holds when both $\mathcal{H}_{K}\cap\mathcal{H}_{u_{0}}\neq\emptyset$ and $\mathcal{H}_{K}\cap\mathcal{H}_{v_{0}}\neq\emptyset$.
Finally, suppose that $K$ intersects both $C_{L}$ and $C_{R}$. Since $(C,\mathcal{K})$ is $axax$-free and $K_{0}$ contains $u_{0}$ and $v_{0}$, it follows by Proposition 19, that $K$ contains either $u_{0}$ or $v_{0}$. Therefore, either $H_{1}$ or $H^{\prime}_{s}$ is in $\mathcal{H}_{K}$. Now, the fact that $\mathcal{H}_{K}$ induces a connected subgraph in $Q$ is identical to the case when $K$ has runs only in $C_{L}$ s.t. $\mathcal{H}_{K}\cap\mathcal{H}_{u_{0}}\neq\emptyset$. Since $K$ was chosen arbitrarily, $Q$ is a support.
Now, we are ready to prove the general case. For the proof of the general case, we require a few technical tools. For a cycle system $(C,\mathcal{H},\mathcal{K})$, let $H_{0}=\arg\min_{H\in\mathcal{H}}\ell(H)$. Let $d_{H_{0}}=\{u_{0},v_{0}\}$ denote the chord of $H_{0}$ attaining the minimum. The two derived cycle systems $(C,\mathcal{H}_{L},\mathcal{K})$ and $(C,\mathcal{H}_{R},\mathcal{K})$ corresponding to $H_{0}$ are defined as follows:
Let $\mathcal{H}_{R}=\{H\cap\mathrm{arc}[u_{0},v_{0}]:H\in\mathcal{H}\setminus H_{0}\}\cup H_{0}^{\prime}$, where $H_{0}^{\prime}$ corresponds to $H_{0}$, and spans all the vertices of the complementary $\mathrm{arc}[v_{0},u_{0}]$. We construct $\mathcal{H}_{L}$ in two steps. Set $\mathcal{H}^{\prime}_{L}=\{H\cap\mathrm{arc}[v_{0},u_{0}]:H\in\mathcal{H}\setminus H_{0}\}$. Let $H\in\mathcal{H}$ be s.t. $H\cap\mathrm{arc}(u_{0},v_{0})\neq\emptyset$. Since $H_{0}\cap\mathrm{arc}(u_{0},v_{0})=\emptyset$ with $u_{0},v_{0}\in H_{0}$, and $(C,\mathcal{H})$ is $axax$-free, it follows from Proposition 19 that in $\mathrm{arc}[v_{0},u_{0}]$, any run of $H$ must be contained in a run of $H_{0}$. Therefore, we remove all such subgraphs from $\mathcal{H}^{\prime}_{L}$. We set $\mathcal{H}_{L}=\mathcal{H}^{\prime}_{L}\setminus\{H\in\mathcal{H}^{\prime}_{L}:H\cap\mathrm{arc}(u_{0},v_{0})\neq\emptyset\}\cup H_{0}^{\prime\prime}$, where $H_{0}^{\prime\prime}=H_{0}\cup\mathrm{arc}[u_{0},v_{0}]$ corresponds to the subgraph $H_{0}$. That is, $H_{0}^{\prime\prime}$ extends from $u_{0}$ and $v_{0}$ to contain the complementary arc $\mathrm{arc}[u_{0},v_{0}]$. This defines the derived cycle systems $(C,\mathcal{H}_{L},\mathcal{K})$ and $(C,\mathcal{H}_{R},\mathcal{K})$ corresponding to the subgraph $H_{0}$. Note that each subgraph $H\in\mathcal{H}\setminus\{H_{0}\}$ has a run in exactly one of $\mathcal{H}_{L}$ or $\mathcal{H}_{R}$. Moreover, if $\mathcal{H}$ is containment-free, and $(C,\mathcal{H})$ is $axax$-free, then it follows that $\mathcal{H}_{R}$ is strict-containment free. Figure 5 shows a partition of $(C,\mathcal{H},\mathcal{K})$ into the derived systems.
((a)) $(C,\mathcal{H})$
((b)) $(C,\mathcal{H}_{L})$
((c)) $(C,\mathcal{H}_{R})$
Figure 5: Collections $\mathcal{H}_{L}$ and $\mathcal{H}_{R}$ of derived systems. $\mathcal{K}$ subgraphs are not shown in the figure.
Proposition 21.
Let $(C,\mathcal{H},\mathcal{K})$ be a strong $axax$-free system. Let $H_{0}=\arg\min_{H\in\mathcal{H}}\ell(H)$ and that $d_{H_{0}}=\{u_{0},v_{0}\}$, where $d_{H_{0}}$ is the chord of $H_{0}$ attaining the minimum. Then, the two derived cycle systems $(C,\mathcal{H}_{L},\mathcal{K})$ and $(C,\mathcal{H}_{R},\mathcal{K})$ corresponding to $H_{0}$ are also strong $axax$-free.
Proof.
Clearly, the cycle systems $(C,\mathcal{H}_{L},\mathcal{K})$ and $(C,\mathcal{H}_{R},\mathcal{K})$ are $axax$-free. We show that they satisfy the intersection property. By the choice of $H_{0}$, in the derived system $(C,\mathcal{H}_{R},\mathcal{K})$, each $H\in\mathcal{H}_{R}$ induces a single run on $C$. Therefore, it is trivially strong $axax$-free. Consider the derived system $(C,\mathcal{H}_{L},\mathcal{K})$. Let $H\in\mathcal{H}_{L}$ and $K\in\mathcal{K}$ be such that there are vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order on $C$ with $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$ such that $H\cap K=\emptyset$. First, suppose that $H=H_{0}^{\prime\prime}$. If $K$ has a vertex in $\mathrm{arc}[u_{0},v_{0}]$, then $H\cap K\neq\emptyset$; a contradiction to our assumption. If $K$ does not have a vertex in $\mathrm{arc}[u_{0},v_{0}]$, then it follows that $H_{0}$ and $K$ does not satisfy the intersection property in $C$ since $H_{0}$ and $H_{0}^{\prime\prime}$ coincide on $\mathrm{arc}[v_{0},u_{0}]$; again a contradiction to the fact that $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free.
So, suppose $H\neq H_{0}^{\prime\prime}$. By construction of $\mathcal{H}_{L}$, $H\cap\mathrm{arc}(u_{0},v_{0})=\emptyset$. This implies $H$ and $K$ do not satisfy the intersection property in $(C,\mathcal{H},\mathcal{K})$, contradicting the assumption that it is strong $axax$-free. ∎
Now we prove the result for the general setting i.e., when subgraphs in $\mathcal{H}\cup\mathcal{K}$ induce more than one run on $C$.
See 14
Proof.
By Lemma 5, we assume wlog that $\mathcal{H}$ is containment-free. We prove the result by induction on $N=N(C,\mathcal{H})=\sum_{H\in\mathcal{H}}(n_{H}-1)\geq 0$. If $N=0$, then each $H\in\mathcal{H}$ induces a single run on $C$, and by Lemma 20, we are done. So, we can assume that $N\geq 1$ i.e., $n_{H}\geq 2$ for some $H\in\mathcal{H}$. Suppose the theorem holds for all strong $axax$-free systems $(C,\mathcal{H}^{\prime},\mathcal{K})$ with $N(C,\mathcal{H}^{\prime})<N$.
Let $(C,\mathcal{H,K})$ be a strong $axax$-free system with $N(C,\mathcal{H})=N$. Let $H_{0}=\arg\min_{H\in\mathcal{H}}\ell(d_{H})$ with $d_{H_{0}}=\{u_{0},v_{0}\}$, and such that $H_{0}\cap\mathrm{arc}(u_{0},v_{0})=\emptyset$.
Consider the derived cycle systems $(C,\mathcal{H}_{L},\mathcal{K})$ and $(C,\mathcal{H}_{R},\mathcal{K})$. By Proposition 21, both derived cycle systems are strong $axax$-free.
In the cycle system $(C,\mathcal{H}_{R},\mathcal{K})$, each subgraph in $\mathcal{H}_{R}$ induces a single run and $\mathcal{H}_{R}$ is strict-containment free. By Lemma 20, there is an outerplanar support $Q_{R}$ whose outer cycle is $C(\mathcal{H}_{R})$. By the choice of $H_{0}$, $N(C,\mathcal{H}_{L})<N$, and by Proposition 21, $(C,\mathcal{H}_{L},\mathcal{K})$ is a strong $axax$-free system. By the inductive hypothesis therefore, $(C,\mathcal{H}_{L},\mathcal{K})$ admits an outerplanar support $Q_{L}$. By construction of $\mathcal{H}_{L}$ and $\mathcal{H}_{R}$, each $H\in\mathcal{H}\setminus\{H_{0}\}$ has its representative vertices in exactly one of $Q_{L}$ or $Q_{R}$. Let $Q$ be the graph obtained by identifying $H^{\prime\prime}_{0}$ in $Q_{L}$ and $H^{\prime}_{0}$ in $Q_{R}$. The graph $Q$ is clearly outerplanar. It remains to show that $Q$ is an intersection support for $(C,\mathcal{H},\mathcal{K})$.
Let $K\in\mathcal{K}$ be arbitrary. We show that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. We consider three cases for the runs of $K$. First, suppose that all runs of $K$ lie in $\mathrm{arc}[u_{0},v_{0}]$. Then all subgraphs in $\mathcal{H}$ intersecting $K$ have a representative in $\mathcal{H}_{R}$. The fact that $\mathcal{H}_{K}$ is connected in $Q$ follows from the fact that $Q_{R}$ is an intersection support for $(C,\mathcal{H}_{R},\mathcal{K})$.
Now suppose that all runs of $K$ lie entirely in $\mathrm{arc}[v_{0},u_{0}]$. If each $H\in\mathcal{H}_{K}$ has its representative in $\mathcal{H}_{L}$, then the fact that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$ follows from the fact that $Q_{L}$ is an intersection support for $(C,\mathcal{H}_{L},\mathcal{K})$. Otherwise, if an $H\in\mathcal{H}_{K}$ does not have its representative in $\mathcal{H}_{L}$, then by construction of $\mathcal{H}_{L}$, $H\cap\mathrm{arc}(u_{0},v_{0})\neq\emptyset$. By Proposition 19 therefore, $H$ contains $u_{0}$ or $v_{0}$ and each run of $H$ in $\mathrm{arc}[v_{0},u_{0}]$ is contained in a run of $H_{0}$. It follows that $K\cap H_{0}\neq\emptyset$ and thus $K\cap H^{\prime\prime}_{0}\neq\emptyset$. Recall that the subgraph in $\mathcal{H}$ containing $u_{0}$ or those containing $v_{0}$ appear consecutively in the outer cycle $C(\mathcal{H}_{R})$ of $Q_{R}$, and $K$ intersects a prefix of these sequences of subgraphs. Since $K$ intersects $H_{0}$ which contains both $u_{0}$ and $v_{0}$, $\mathcal{H}_{K}$ induces a connected subgraph of $Q$.
Finally, let $K$ intersect both $\mathrm{arc}[u_{0},v_{0}]$ and $\mathrm{arc}[v_{0},u_{0}]$. Note that in this case $K$ intersects $H_{0}$ (and hence with $H^{\prime}_{0}$ and $H^{\prime\prime}_{0}$) since $u_{0},v_{0}\in H_{0}$ and $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. Each $H\in\mathcal{H}\setminus\{H_{0}\}$ has a representative vertex in one of $Q_{L}$ or $Q_{R}$ which are supports for the derived cycle systems with $H_{0}$ in both $Q_{L}$ and $Q_{R}$. It follows that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. Since $K$ was chosen arbitrarily, $Q$ is a support.
5.3 Implementation
In this section, we show a polynomial running time of our algorithm to construct an outerplanar support. First, we give a short argument that $|\mathcal{H}|$ and $|\mathcal{K}|$ are polynomially bounded. Possibly the bound here is not sharp, but is sufficient for our purposes since we only want to establish a polynomial running time. We use the results of [ackerman2020coloring] and [RR18].
Theorem 22.
Let $(G,\mathcal{H},\mathcal{K})$ be an outerplanar non-piercing graph system. Then, $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$, where $n=|V(G)|$.
Proof.
Let $(C,\mathcal{H},\mathcal{K})$ be the resulting cycle system. By Lemma 13, $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free, and this implies that $(C,\mathcal{H})$ and $(C,\mathcal{K})$ are $abab$-free.
Ackerman et al. [ackerman2020coloring] showed that $abab$-free hypergraphs are exactly those that can be represented by a set of pseudodisks [^5] A family $\mathcal{D}$ of compact regions in the plane s.t. the boundary of each region $D\in\mathcal{D}$ is a simple Jordan curve and for any two regions $D,D^{\prime}$ their boundaries intersect either 00 or 2 times is called a set of pseudodisks. that contain a common point. Hence, the hypergraphs defined by $\mathcal{H}$ can be represented s.t. the vertices of $C$ are points in the plane and the hyperedges in $\mathcal{H}$ are pseudodisks containing the origin, and each hyperedge is defined by the set of points of $C$ in the pseudodisk.
We show that the VC-dimension of the set system $(C,\mathcal{H})$ is at most 4. The results of Raman and Ray [RR18] imply that for any set $P$ of points and any set $\mathcal{D}$ of pseudodisks, there is a planar support $Q$ on $P$, i.e., a planar graph on $P$ s.t. the points in $D$ induce a connected subgraph of $Q$ for each $D\in\mathcal{D}$. If a set of 5 points can be shattered, then for each pair of points, there is a hyperedge containing that pair. But, this implies that support for these 5 points is $K_{5}$, contradicting the fact that it is planar.
Hence, the VC-dimension of the set system $(C,\mathcal{H})$ is at most 4, and this implies, by the Sauer-Shelah lemma (See Chapter 10 in [Mat02LecDiscGeom]) that $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$.
Theorem 23.
If $(G,\mathcal{H},\mathcal{K})$ is an outerplanar non-piercing intersection system, then an intersection support can be computed in time $O(n^{6})$ where $n=|V(G)|$.
Proof.
Let $(C,\mathcal{H},\mathcal{K})$ be the resulting strong $axax$-free cycle system. Then $|C|=O(n)$, and by Theorem 22, $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$. We only need to show that our algorithm in Theorem 14 runs in time $O(n^{6})$.
If each $X\in\mathcal{H}\cup\mathcal{K}$ induces a single run on $C$, then by Lemma 18, $C(\mathcal{H})$ is the desired support. To construct $C(\mathcal{H})$, we walk along $C$ and at each vertex $v$, add the subgraphs $H\in\mathcal{H}$ to $C(\mathcal{H})$ s.t. $s(H)=v$. For each vertex, we order the subgraphs in increasing order of $t(H)$, which can be computed by sorting the subgraphs $H$ with $s(H)=v$. Thus, $C(\mathcal{H})$ can be computed in time $O(|C|+|\mathcal{H}|\log|\mathcal{H}|)=O(n^{4}\log{n})$.
For the case when $\mathcal{H}$ consists of single runs and a subgraph in $\mathcal{K}$ can have multiple runs, finding a chord $d$ of smallest length can be done by storing the subgraphs in $\mathcal{K}$ in a heap ordered by $\ell(K)$. The time taken to partition the problem into two sub-problems is $O(|C|\max\{|\mathcal{H}|,|\mathcal{K}|\})$ as we need to go through each subgraph in $\mathcal{H}$ and $\mathcal{K}$, and split the runs into the two sub-problems. We add at most $|C|-3$ chords in $C$ since the resulting support is outerplanar. Hence, the total running time in this case is $O(|C|^{2}|\max\{|\mathcal{H}|,|\mathcal{K}|\})=O(n^{6})$.
In the general case, a subgraph in $\mathcal{H}$ can have multiple runs. To compute an $H\in\mathcal{H}$ minimizing $\ell(H)$, we can store the subgraphs in $\mathcal{H}$ in a heap ordered by $\ell(H)$. We can split the problem into two sub-problems in $O(|C|\max\{|\mathcal{H}|,|\mathcal{K}|\})$ time. Since we add at most $|C|-3$ chords, the overall running time is bounded above by $O(|C|^{2}\max\{|\mathcal{H}|,|\mathcal{K}|\})$. Hence, the overall running time of the algorithm is $O(n^{6})$.
Outerplanar graphs have treewidth at most $2$ and the above result shows that if the subgraphs are non-piercing, then the VC dimension of the set system defined by $(C,\mathcal{H})$ is at most $4$. We extend this result in Section 6 to graphs of bounded treewidth and we show that for a non-piercing graph system $(G,\mathcal{H})$ of treewidth $t$, the VC-dimension of set system defined by $(V(G),\mathcal{H})$ is at most $33t+$ 3.
5.4 Primal and Dual Supports: Revisited
The existence of an intersection support also implies a primal and a dual support. However, we emphasize here that for a primal and a dual support, we require strictly weaker conditions. In particular, for a primal support, we require the cycle system $(C,\mathcal{H})$ to be $abab$-free and not necessarily $axax$-free, and for a dual support, we require $(C,\mathcal{H})$ to be $axax$-free. The result for a primal support directly follows from the following result of Raman and Singh [raman2025supportsgraphsboundedgenus].
Lemma 24 ([raman2025supportsgraphsboundedgenus]).
Let $(C,\mathcal{H})$ be an $abab$-free cycle system. Then, we can add a set $D$ of non-intersecting chords in $C$ such that each $H\in\mathcal{H}$ induces a connected subgraph of $C\cup D$. Further, the set $D$ of non-intersecting chords to add can be computed in time $O(mn^{4})$ where $n=|C|$, and $m=|\mathcal{H}|$.
Theorem 25.
Let $(C,\mathcal{H})$ be an $abab$-free cycle system and $c:V(C)\to\{\mathbf{r},\mathbf{b}\}$ be a 2-coloring of the vertices of $C$. Then there is a primal support for $(C,\mathcal{H})$ that is an outerplanar graph.
Proof.
For each $v\in V(C)$ such that $c(v)=\mathbf{r}$, we delete it and make its neighbours adjacent. Let $C^{\prime}$ be the resulting cycle obtained form $C$, and =$\mathcal{H}^{\prime}=\{H\cap C^{\prime}:H\in\mathcal{H}\}$. Then $(C^{\prime},\mathcal{H}^{\prime})$ is an $abab$-free cycle system. By Lemma 24, there is a set $D$ of non-intersecting chord added to $C^{\prime}$ such that each subgraph in $\mathcal{H}^{\prime}$ induces a connected subgraph of $C^{\prime}\cup D$. Thus $C^{\prime}\cup D$ is an outerplanar primal support for $(C^{\prime},\mathcal{H}^{\prime})$. Note that this is also a primal support for $(C,\mathcal{H})$.
For the dual setting, one may likewise hope that the $abab$-free condition is sufficient to obtain a dual support. Unfortunately, that is not the case as the following example shows: Let $G$ be the graph on $\{1,2,\ldots 6\}$ as shown in Figure 6. Let $\mathcal{H}$ be subgraphs $H_{1}=\{1,2,3\}$, $H_{2}=\{3,4,5\}$, $H_{3}=\{5,6,1\}$ and $H_{4}=\{2,4,6\}$. It can be checked that $\mathcal{H}$ is $abab$-free. The dual support for $(G,\mathcal{H})$ is $K_{4}$, a complete graph on 4 vertices, which is not outerplanar.
The reason why $(G,\mathcal{H})$ does not admit a dual outerplanar support is that the induced system $(C,\mathcal{H})$ is not $axax$-free - $H_{4}$ forms an $axax$-pair with each of $H_{1},H_{2}$ and $H_{3}$. However, if $axax$-freeness is a sufficient condition for $(C,\mathcal{H})$ to exhibit an outerplanar dual support and it follows from Theorem 14.
Theorem 26.
If $(C,\mathcal{H})$ is $axax$-free, then it admits an outerplanar dual support.
Proof.
Let $\mathcal{K}=\{\{v\}:v\in V(C)\}$ be a collection of subgraphs each consisting of a single vertex. Then $(C,\mathcal{K})$ is $axax$-free. Therefore, the intersection system $(C,\mathcal{H},\mathcal{K})$ is $axax$-free. Note that the strong $axax$-free property is trivially satisfied by $(C,\mathcal{H},\mathcal{K})$. By Theorem 14, it has an intersection support for $Q$, which is also a dual support for $(C,\mathcal{H})$ since each $v\in V(C)$ corresponds to a subgraph in $K\in\mathcal{K}$.
Figure 6: $abab$-free cycle system whose only dual support is $K_{4}$. The induced subgraphs are shown by four different colors on edges.
Again, we might hope that for an outerplanar intersection system $(G,\mathcal{H},\mathcal{K})$, if it satisfies the $axax$-free property, then like the primal and dual, we can obtain an outerplanar intersection support. However, the following example shows that in order to construct an intersection support, one needs the strong $axax$-free property: Let $C=(1,2,\ldots 7,1)$ be a cycle as shown in Figure 7(a). Let $\mathcal{H}$ consist of induced subgraphs $H_{1}=\{1,5\},H_{2}=\{2\},H_{3}=\{3,4\}$ and $H_{4}=\{6,7\}$, and $\mathcal{K}$ consist of induced subgraphs $K_{1}=\{1,2\},K_{2}=\{2,3\},K_{3}=\{4,6\},K_{4}=\{1,7\},K_{5}=\{2,7\}$ and $K_{6}=\{4,5\}$ of $C$. It is easy to check that both $(C,\mathcal{H})$ and $(C,\mathcal{K})$ are $axax$-free. However, for each $K_{i}$, $i=1,\ldots,6$, there is exactly one pair $H_{k},H_{\ell}$, $k\neq\ell\in\{1,\ldots,4\}$ s.t. $H_{k}$ and $H_{\ell}$ intersect $K_{i}$, and hence, the intersection support is a complete graph on $4$ vertices (see Figure 7(b)), which is not outerplanar.
Observe that vertices $1,4,5,6$ are in cyclic sequence with $1,5\in H_{1}$ and $4,6\in K_{3}$ such that $H_{1}\cap K_{3}=\emptyset$. Similarly, vertices $1,2,5,7$ appear in cyclic sequence with $1,5\in H_{1}$ and $2,7\in K_{5}$ such that $H_{1}\cap K_{5}=\emptyset$. Hence, the graph system $(C,\mathcal{H},\mathcal{K})$ does not satisfy the strong $axax$-free property.
((a)) Alternate sequences of $H_{1}$ with $K_{3}$ and $K_{5}$.
((b)) Non-outerplanar support. Edge $e_{i}$ corresponds to $K_{i}$ for $i=1\ldots 6$.
Figure 7: An example of an $axax$-free, but not strong $axax$-free system $(C,\mathcal{H,K})$ that does not admit an outerplanar support.
6 Support of Bounded Treewidth
In this section, we show that if $(G,\mathcal{H,K})$ is a non-piercing system of treewidth $t$, then there exists an intersection support of treewidth $2^{O(2^{t})}$. In order to show this, we construct a primal and dual supports of treewidth $O(2^{t})$ and $O(2^{4t})$ respectively, and use these to construct an intersection support. Further, our construction is FPT in $\textsc{tw}(G)$. First, we collect some tools required for the construction of supports.
For a graph system $(G,\mathcal{H})$, let $(T,\mathcal{B})$ denote a tree decomposition of $G$ of width $t$. Let $CC(G)$ denote the chordal completion of $G$, i.e., we add edges so that all vertices in a bag are pairwise adjacent. Clearly, a chordal completion does not increase the width of the tree decomposition. Thus if $(T,\mathcal{B})$ is of width $\textsc{tw}(G)$, then $\textsc{tw}(CC(G))=\textsc{tw}(G)$. Moreover, if $\mathcal{H}$ is a collection of non-piercing subgraphs of $G$, they remain non-piercing when induced on $CC(G)$ since the underlying hypergraph does not change. Further, it is straightforward that computing a support w.r.t. $CC(G)$ is equivalent to computing a support w.r.t. $G$. Hence, we assume throughout this chapter that, $G$ is a chordal graph [^6] A graph is chordal, if there is no induced cycle of length at least 4..
We use the following notations: We assume throughout this section that $T$ be a binary tree rooted at node $\rho$. For any node $x$ of $T$, we use $T_{x}$ to denote the sub-tree rooted at $x$. Let $\mathcal{B}_{x}$ denote the set of bags at the nodes in $T_{x}$. The subgraph $G_{x}=\cup_{z\in T_{x}}B_{z}$, of $G$ is the graph induced on the vertices corresponding to the bags in $\mathcal{B}_{x}$. Note that $(T_{x},\mathcal{B}_{x})$ induces a tree decomposition of $G_{x}$. We use $G_{-x}$ to denote the subgraph induced on $V(G)\setminus V(G_{x})$.
Let $\mathcal{H}$ be a collection of subgraphs of $G$. For $H\in\mathcal{H}$, we let $H|_{x}=G_{x}\cap H$ and $\mathcal{H}|_{x}=\{H|_{x}:H\in\mathcal{H}\}$. Similarly, $H|_{-x}=H\cap G_{-x}$ and $\mathcal{H}|_{-x}=\{H|_{-x}:H\in\mathcal{H}\}$. Then, $(G_{x},\mathcal{H}|_{x})$ and $(G_{-x},\mathcal{H}_{-x})$ denote the two induced graph systems on $G_{x}$ and $G_{-x}$ respectively.
For an edge $\{x,y\}$ of $T$, we use $A_{xy}$ to denote the adhesion set $B_{x}\cap B_{y}$. For $A\subseteq V(G)$, let $\mathcal{H}_{A}=\{H\in\mathcal{H}:A\cap H\neq\emptyset\}$, and for $S\subseteq A$, $\mathcal{H}^{=}_{S}=\{H\in\mathcal{H}_{A}:H\cap A=S\}$.
The next two lemmas follow from the fact that the subgraphs considered are non-piercing. For two sets $A$ and $B$ on the same ground set, we say that $A$ and $B$ properly intersect if $A\setminus B\neq\emptyset$ and $B\setminus A\neq\emptyset$.
Lemma 27.
Let $(G,\mathcal{H})$ be a non-piercing graph system with a tree decomposition $(T,\mathcal{B})$ of $G$. Let $e=\{x,y\}\in E(T)$ be an edge in $T$ where $x$ is a child of $y$. Then for $H,H^{\prime}\in\mathcal{H}_{A_{xy}}$, the following holds: $(i)$ If $A_{xy}\cap H=A_{xy}\cap H^{\prime}$ and $H|_{x}\subset H^{\prime}|_{x}$, then $H^{\prime}|_{-x}\subseteq H|_{-x}$. $(ii)$ If $A_{xy}\cap H=A_{xy}\cap H^{\prime}$, and $H|_{x}$ and $H^{\prime}|_{x}$ properly intersect, then $H|_{-x}=H^{\prime}|_{-x}$, and $(iii)$ If $H\cap A_{xy}\subset H^{\prime}\cap A_{xy}$, and $H|_{x}$ and $H^{\prime}|_{x}$ properly intersect, then $H|_{-x}\subseteq H^{\prime}|_{-x}$.
Proof.
The proofs follow immediately from the fact that $\mathcal{H}$ are non-piercing and the adhesion set corresponding to any edge in $E(T)$ is a separator in $G$.
Lemma 28.
Let $(G,\mathcal{H})$ be a non-piercing graph system with a tree decomposition $(T,\mathcal{B})$ of $G$ For any node $x$ of $G$, the graph system $(G_{x},\mathcal{H}_{x})$ is non-piercing and $(T_{x},\mathcal{B}_{x})$ is a tree decomposition of $G_{x}$.
Proof.
Since $G$ is chordal, each bag induces a complete subgraph of $G$. Hence, each $H_{x}\in\mathcal{H}_{x}$ is connected. If $x$ is not a root of $T$, let $y$ be the parent of $x$. Then, for $H_{x},H^{\prime}_{x}\in\mathcal{H}_{x}$, the subgraph $H_{x}\setminus H^{\prime}_{x}$ of $G_{x}$ is connected since $H\setminus H^{\prime}$ is connected and $A_{xy}$ is a separator in $G$. Therefore, $(G_{x},\mathcal{H}_{x})$ is a non-piercing graph system. Finally, that $(T_{x},\mathcal{B}_{x})$ is a tree decomposition, is straightforward.
6.1 Primal Support
In this section, we show that a bounded treewidth non-piercing system $(G,\mathcal{H})$ with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ admits a primal support $Q$ s.t. $\textsc{tw}(Q)\leq 2^{\textsc{tw}(G)+2}+\textsc{tw}(G)$. The proof is algorithmic and yields a polynomial time algorithm if $\textsc{tw}(G)$ is bounded, i.e., an FPT-algorithm parameterized by $\textsc{tw}(G)$.
An easy tree decomposition of $G$ is a tree decomposition s.t. for each adhesion set $A$ of the tree decomposition and each subgraph $H$ intersecting $A$, it does so at a blue vertex, i.e., for any adhesion set $A$, if $H\cap A\neq\emptyset$, then $H\cap\mathbf{b}(A)\neq\emptyset$. If $G$ has an easy tree decomposition of treewidth $t$, then it is straightforward to construct a primal support $Q$ s.t. $\textsc{tw}(Q)\leq t$. Moreover, in this case, we only require the subgraphs in $\mathcal{H}$ to be connected.
Lemma 29.
Let $(G,\mathcal{H})$ be a graph system with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ s.t. each $H\in\mathcal{H}$ is a connected subgraph of $G$. If there is an easy tree decomposition $(T,\mathcal{B})$ of width $t$, then there is a primal support $Q$ on $\mathbf{b}(V)$ of treewidth at most $t$.
Proof.
Let $(T,\mathcal{B}^{\prime})$ be the tree decomposition on $\mathbf{b}(V)$ derived from $(T,\mathcal{B})$, where $B^{\prime}=B\cap\mathbf{b}(B)$ is the bag in $\mathcal{B}^{\prime}$ corresponding to the bag $B\in\mathcal{B}$. For each $B^{\prime}\in\mathcal{B}^{\prime}$, the vertices of $B^{\prime}$ induce a clique since we work with the chordal completion of $G$. Then, $(T,\mathcal{B}^{\prime})$ is the tree decomposition of a graph $Q$.
To show that $Q$ is a support, consider an $H\in\mathcal{H}$. Since $H$ is a connected subgraph in $G$, the vertices in $V(H)$ lie in bags corresponding to a connected sub-tree of $T$. Further, as $(T,\mathcal{B})$ is an easy tree decomposition, $\mathbf{b}(V(H))$ lie in bags of $\mathcal{B}^{\prime}$ corresponding to a connected sub-tree of $T$. Since each $B^{\prime}\in\mathcal{B}^{\prime}$ induces a complete subgraph of $Q$, it implies $V(H)$ induces a connected subgraph of $Q$.
In the proof below, we use the following notation: Let $\rho$ be the parent of $x$ in the tree decomposition $(T,\mathcal{B})$. For an adhesion set $A_{x\rho}$ and $S\subseteq A_{x\rho}$, for notational convenience, we assume that each $H\in\mathcal{H}^{=}_{S}$ is s.t. $H|_{x}\cap\mathbf{b}(G_{x})\neq\emptyset$ and $H|_{-x}\cap\mathbf{b}(G_{-x})\neq\emptyset$. Let $\mathcal{M}_{S}\subseteq\mathcal{H}^{=}_{S}$ denote the set of subgraphs $H$ such that $H_{x}$ is minimal in $G_{x}$ in the containment order $\preceq$, i.e., for $H,H^{\prime}\in\mathcal{H}^{=}_{S}$, $H|_{x}\preceq H^{\prime}|_{x}\Leftrightarrow H|_{x}\subseteq H^{\prime}|_{x}$. We use $(G_{x},\preceq)$ to denote this containment order. We use minimal elements in $\mathcal{M}_{S}$ to construct an easy tree decomposition at the cost of an increase in the width of the tree decomposition.
Lemma 30.
Let $(G,\mathcal{H})$ be a non-piercing graph system with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$. A tree decomposition $(T,\mathcal{B})$ of $G$ of width $t$ can be transformed into a tree decomposition $(T,\mathcal{B}^{\prime})$ of width at most $2^{t+2}+t$.
Proof.
If $(T,\mathcal{B})$ is an easy tree decomposition, we are done. Otherwise, we modify $(T,\mathcal{B})$ to an easy tree decomposition $(T,\mathcal{B}^{\prime})$ by adding additional blue vertices to the bags in $\mathcal{B}$. We assume without loss of generality that $(T,\mathcal{B})$ is a binary tree rooted at a node $\rho$.
We prove by induction on the height of $T$. If $T$ has height 0, then $T$ consists of a single node $\rho$ and hence for each $H\in\mathcal{H}$, $H\cap\mathbf{b}(B_{\rho})\neq\emptyset$. Otherwise, let $x$ and $y$ be children of $\rho$. By Lemma 28, the graph system $(G_{x},\mathcal{H}_{x})$ is non-piercing and $(T_{x},\mathcal{B}_{x})$ is a tree decompositions of $G_{x}$. By the inductive hypothesis, there is an easy tree decomposition $(T_{x},\mathcal{B}^{\prime}_{x})$ of width at most $2^{t+2}+t$. Analogously, there is an easy tree decomposition $(T_{y},\mathcal{B}^{\prime}_{y})$ of width at most $2^{t+2}+t$ for $(G_{y},\mathcal{H}_{y})$.
For each $S\subseteq\mathbf{r}(A_{x\rho})$, consider an $H\in\mathcal{M}_{S}$. Since $H\cap B^{\prime}_{x}\neq\emptyset$ and $(T,\mathcal{B}^{\prime}_{x})$ is an easy tree decomposition, $H\cap\mathbf{b}(B^{\prime}_{x})\neq\emptyset$. Choose a $b\in H\cap\mathbf{b}(B^{\prime}_{x})$ and add it to $B_{\rho}$. The tree decomposition remains a valid tree decomposition as the bags containing $b$ correspond to a connected subset of nodes of $T$. Similarly, for each $S\subseteq\mathbf{r}(A_{y\rho})$, choose an $H^{\prime}\in\mathcal{M}_{S}$ and a vertex $b^{\prime}\in H^{\prime}\cap\mathbf{b}(B^{\prime}_{y})$ and add $b^{\prime}$ to $B_{\rho}$. Let $B^{\prime}_{\rho}$ denote the bag at $\rho$ after all subsets of $A_{x\rho}$ and $A_{y\rho}$ have been processed. Since $|A_{x\rho}|,|A_{y\rho}|\leq t+1$ and $|B_{\rho}|\leq t+1$, we have $|B^{\prime}_{\rho}|\leq 2\cdot 2^{t+1}+(t+1)$. Therefore, width of the tree decomposition $(T,\mathcal{B}^{\prime})$ is at most $2^{t+2}+t$.
We claim that $(T,\mathcal{B}^{\prime})$ is an easy tree decomposition. Suppose not. Let $H\in\mathcal{H}$ s.t. $H\cap B_{\rho}\neq\emptyset$ and $H\cap\mathbf{b}(B^{\prime}_{\rho})=\emptyset$. Then, $H\cap A_{x\rho}\neq\emptyset\neq H\cap A_{y\rho}$ by assumption. Let $H\cap A_{x\rho}=S$ and $H\cap A_{y\rho}=S^{\prime}$. Let $H^{\prime}$ be the minimal subgraph in $\mathcal{M}_{S}$ whose blue vertex was added to $B^{\prime}_{\rho}$ and let $H^{\prime\prime}$ be the minimal subgraph in $\mathcal{M}_{S^{\prime}}$ whose blue vertex was added to $B^{\prime}_{\rho}$. Then, $H$ and $H^{\prime}$ are incomparable in $(G_{x},\preceq)$. By part $(ii)$ of Lemma 27, this implies $H$ and $H^{\prime}$ are identical in $G_{-x}$. Similarly, $H$ and $H^{\prime\prime}$ are incomparable in $(G_{y},\preceq)$ and by Lemma 27 $H$ and $H^{\prime\prime}$ are identical in $G_{-y}$. It follows that $H^{\prime}|_{x}$ and $H^{\prime\prime}|_{x}$ are incomparable in $G_{x}$, and $H^{\prime}|_{y}$ and $H^{\prime\prime}|_{y}$ are incomparable in $G_{y}$. Also, note that $H^{\prime\prime}\cap A_{x\rho}=S=H^{\prime}\cap A_{x\rho}$. But this implies $H^{\prime}\setminus H^{\prime\prime}$ has at least two connected components - one in $G_{x}$ and one in $G_{y}$ contradicting the assumption that the subgraphs in $\mathcal{H}$ are non-piercing. Hence, $(T,\mathcal{B}^{\prime})$ is an easy tree decomposition.
See 7
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$ where $T$ is a binary rooted tree. If $(T,\mathcal{B})$ is an easy tree decomposition, we are done by Lemma 29. Otherwise, we transform it into an easy tree decomposition of width at most $2^{t+2}+t$ by Lemma 30. We then obtain a primal support $Q$ of the same treewidth by Lemma 29.
6.2 Dual Support
We now construct a dual support. For a non-piercing system $(G,\mathcal{H})$ of treewidth $t$, we construct a dual support of treewidth $O(2^{4t})$. By Lemma 5, we assume that there are no containments, i.e., there are no two subgraphs $H,H^{\prime}\in\mathcal{H}$ such that $H\subseteq H^{\prime}$. As in the primal setting, we first show how we can obtain a dual support for a simple case. A support for the general case is obtained by reducing it to the simple case. In the following, since we construct a graph on $\mathcal{H}$, we abuse notation and also use $H$ to denote the vertex in the dual support corresponding to $H\in\mathcal{H}$.
For a graph system $(G,\mathcal{H})$, a tree decomposition $(T,\mathcal{B})$ of $G$ is said to be k-sparse with respect to $\mathcal{H}$ if for each bag $B\in\mathcal{B}$, at most $k$ subgraphs in $\mathcal{H}$ intersect $B$. If $(G,\mathcal{H})$ admits such a tree decomposition, we say that $(G,\mathcal{H})$ is $k$-sparse with respect to $\mathcal{H}$. Note that we do not restrict $\mathcal{H}$ to be non-piercing for $(G,\mathcal{H})$ to be $k$-sparse.
For a connected graph system $(G,\mathcal{H})$ with a tree decomposition $(T,\mathcal{B})$ of width $t$ that is $k$-sparse with respect to $\mathcal{H}$, a dual support can be computed in time $O(k^{2}t|V(G)||\mathcal{H}|)$. The algorithm, $k$-SDS that achieves this is as follows: We construct a tree decomposition $(T,\mathcal{B}^{\prime})$ of a graph $Q^{*}$. For each bag $B\in\mathcal{B}$, we put a vertex $H$ in the corresponding bag $B^{\prime}\in\mathcal{B}^{\prime}$ if $H\cap B\neq\emptyset$, and then put a complete graph on these vertices. $(T,\mathcal{B}^{\prime})$ is the desired tree decomposition of $Q^{*}$ on $\mathcal{H}$. Since each $H\in\mathcal{H}$ is connected and each bag in $\mathcal{B}^{\prime}$ induces a clique, it is clear that $(T,\mathcal{B}^{\prime})$ is a valid tree decomposition. Further, since $(T,\mathcal{B})$ is $k$-sparse, it implies that $(T,\mathcal{B}^{\prime})$ has treewidth at most $k$.
Lemma 31.
For a connected graph system $(G,\mathcal{H})$ and a tree decomposition $(T,\mathcal{B})$ of width $t$ that is $k$-sparse with respect to $\mathcal{H}$, Algorithm $k$-SDS computes a dual support $Q^{*}$ with $\textsc{tw}(Q^{*})\leq k$.
Proof.
For a vertex $v\in V$, consider $\mathcal{H}_{v}$, the set of all subgraphs containing $v$. Since $(T,\mathcal{B})$ is a tree decomposition, at each node $x\in V(T)$ s.t. $v\in B_{x}$, $\mathcal{H}_{v}\subseteq B^{\prime}_{x}$, where $B^{\prime}_{x}$ is the bag at node $x$ in $\mathcal{B}^{\prime}$, i.e., corresponding to $B_{x}$. Since the vertices in $B^{\prime}_{x}$ are pairwise adjacent, it implies $\mathcal{H}_{v}$ induces a connected subgraph of $Q^{*}$. Hence, $Q^{*}$ is a dual support.
For the general setting, we obtain a dual support by sparsifying the input graph system so that it is $k$-sparse for some $k$, and such that a support on the sparsified graph system yields a support for the original graph system. The sparsification yields a set $\mathcal{H}^{\prime}$ of subgraphs that satisfy the following properties: $(i)$ they are in bijective correspondence with $\mathcal{H}$, $(ii)$ each $H^{\prime}\in\mathcal{H}^{\prime}$ corresponding to $H\in\mathcal{H}$ is s.t. $H^{\prime}\subseteq H$. $(iii)$ If $H^{\prime}\subset H$, then there is an $H^{\prime\prime}\in\mathcal{H}$ that pushed out $H$. In this case, $H^{\prime}=H\setminus H^{\prime\prime}$ and there is an edge $e=\{u,v\}\in E(G)$ s.t. $u\in H^{\prime}$ and $v\in H^{\prime\prime}$. We call $e$ a connecting edge between $H^{\prime}$ and $H^{\prime\prime}$.
$\mathcal{H}^{\prime}$ may contain multiple subgraphs spanning the same set of vertices. We denote by $\textrm{unique}(\mathcal{H}^{\prime})$, a subset of $\mathcal{H}^{\prime}$ consisting of one representative from each set of identical subgraphs.
Lemma 32.
Let $(G,\mathcal{H})$ be a non-piercing graph system, and $(T,\mathcal{B})$ a tree decomposition of $G$ of width $t$. Then, there is a connected graph system $(G,\mathcal{H}^{\prime})$ s.t. $(T,\mathcal{B})$ is a $2^{4(t+1)}$-sparse with respect to $\textrm{unique}(\mathcal{H}^{\prime})$. If $Q^{\prime}$ is a dual support for $(G,\textrm{unique}(\mathcal{H}^{\prime}))$, there is a dual support $Q^{*}$ for $(G,\mathcal{H}^{\prime})$ s.t. $\textsc{tw}(Q^{*})=\textsc{tw}(Q^{\prime})$ and $Q^{*}$ is also a dual support for $(G,\mathcal{H})$.
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of width $t$. We assume wlog that $T$ is a binary tree rooted at $\rho$. To obtain $\mathcal{H}^{\prime}$, we do a post-order edge traversal of $T$ and at each edge, we do a pushing.
The pushing operation is as follows: Let $z$ be the parent of a node $x$ in $T$. Having done the pushing on the edges in $T_{x}$, we do the following at $\{x,z\}$: For each $\emptyset\neq S\subseteq A_{xz}$, choose an $H\in\mathcal{M}_{S}$. $H$ is called the pusher for $S$ at $A_{xz}$. For each $H^{\prime}\in\mathcal{H}^{=}_{S}$ s.t. $H^{\prime}|_{x}\setminus H|_{x}\neq\emptyset$, replace $H^{\prime}$ by $H^{\prime}|_{x}\setminus H|_{x}$ in $\mathcal{H}$. We say that $H^{\prime}$ has been pushed by $H$ at $A_{xz}$.
Observe that once a subgraph $H^{\prime}$ is pushed by a pusher $H$ at $A_{xz}$, $H^{\prime}|_{x}\setminus H|_{x}\subseteq G_{x}$, and since pushing is done in a post-order edge traversal of $T$, $H^{\prime}$ is pushed at most once. Further, the subgraphs in $\mathcal{H}$ intersecting $A_{xz}$ were non-piercing before pushing and since $H^{\prime}|_{x}\setminus H|_{x}\neq\emptyset$, by part $(i)$ and $(ii)$ of Lemma 27, $H^{\prime}|_{-x}\subseteq H|_{-x}$. Hence, $H^{\prime}|_{x}\setminus H|_{x}=H^{\prime}\setminus H$. This implies $H^{\prime}|_{x}\setminus H|_{x}$ is a connected subgraph of $G$. It follows that there is a connecting edge $\{u,v\}\in E(G)$ s.t. $u\in H^{\prime}|_{x}\setminus H|_{x}$ and $v\in H|_{x}$ as $H^{\prime}|_{x}\setminus H|_{x}\neq\emptyset$ and $H^{\prime}\cap H\neq\emptyset$.
Let $\mathcal{H}^{\prime}$ be the subgraphs at the end of the algorithm. Note that by construction $\mathcal{H}^{\prime}$ is in bijective correspondence with $\mathcal{H}$. For $S\subseteq A_{xz}$, if a subgraph $H^{\prime}\in\mathcal{H}^{=}_{S}$ was not pushed by a pusher $H$ at $A_{xz}$, then $H^{\prime}|_{x}=H|_{x}$. Therefore, the number of distinct subgraphs of $G_{x}$ in the collection $\mathcal{H}^{\prime}$ intersecting $A_{xz}$ is less than $2^{t+1}$. That is,
$|\textrm{unique}(\mathcal{H}^{\prime}_{A_{xz}}|_{x})|<2^{t+1}$ (1)
The subgraphs in $\textrm{unique}(\mathcal{H}^{\prime})$ intersecting the bag $B_{x}$ can be associated with 4-tuples based on its intersection with $A_{xz}$, with the adhesion sets between $x$ and its children, or with the bag $B_{x}$ itself. From Eqn. (1), it follows that there are at most $2^{4(t+1)}$ distinct subgraphs in $\textrm{unique}(\mathcal{H}^{\prime})$ intersecting $B_{x}$.
By the arguments above, $(G,\textrm{unique}(\mathcal{H}^{\prime}))$ is a connected graph system that is at most $2^{4(t+1)}$-sparse, and hence by Lemma 31, it has a dual support $Q^{\prime}$ of treewidth at most $2^{4(t+1)}$. By Lemma 5, we can extend $Q^{\prime}$ to a support $Q^{*}$ for $(G,\mathcal{H}^{\prime})$ without increasing the treewidth.
We now argue that $Q^{*}$ is a dual support for $(G,\mathcal{H})$. Consider a vertex $v\in V(G)$. The algorithm above ensures that each subgraph is pushed at most once. Suppose $H^{\prime}\in\mathcal{H}_{v}$ was pushed by $H$ so that its modified copy $H^{\prime\prime}=H^{\prime}\setminus H$ does not cover $v$. Then, $H\in\mathcal{H}_{v}$. Let $e=\{a,b\}$ be the connecting edge between $H$ and $H^{\prime\prime}$ such that $a\in H$ and $b\in H^{\prime\prime}$. Since $(T,\mathcal{B})$ is a valid tree-decomposition, there is a bag $B$ containing both $a$ and $b$. Let $H_{1}$ be the unique representative of $H^{\prime\prime}$ in $\textrm{unique}(\mathcal{H}^{\prime})$. Since Algorithm $k$-SDS puts a complete graph on the subgraphs intersecting $B$, it implies $H$ and $H_{1}$ are adjacent in $Q^{\prime}$. In $Q^{*}$, we made $H^{\prime}$ adjacent to $H_{1}$. Hence, $Q^{*}$ is a dual support for $(G,\mathcal{H})$.
See 8
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. If $(T,\mathcal{B})$ is $2^{4(t+1)}$-sparse, then we are done. Otherwise, by Lemma 32, we obtain a dual support $Q^{*}$ for $(G,\mathcal{H})$, of treewidth at most $2^{4(t+1)}$.
6.3 Intersection Support
In this section, we obtain an intersection support of treewidth $2^{O(2^{\textsc{tw}(G)})}$ for a non-piercing intersection system $(G,\mathcal{H},\mathcal{K})$. The construction of the intersection support uses the construction of both the primal and dual support and this leads to the double exponential bound on the treewidth of the intersection support.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$. $(T,\mathcal{B})$ is said to be a $\mathcal{K}$-easy tree decomposition if for each $K\in\mathcal{K}$ and each adhesion set $A$ with $A\cap K\neq\emptyset$, there is an $H\in\mathcal{H}_{K}$ s.t. $H\cap(K\cap A)\neq\emptyset$. Similar to the setting for the dual support, we say that $(T,\mathcal{B})$ is $k$-sparse with respect to $\mathcal{H}$ if for each bag $B\in\mathcal{B}$, there are at most $k$ subgraphs in $\mathcal{H}$ that intersect $B$.
We start by showing that if $(T,\mathcal{B})$ is a $\mathcal{K}$-easy $k$-sparse tree decomposition, then there is an intersection support of treewidth at most $k$. The proof follows along the same lines as the proofs of Lemma 29 and Lemma 31.
Lemma 33.
Let $(G,\mathcal{H,K})$ be a connected intersection system. If $(T,\mathcal{B})$ is a $\mathcal{K}$-easy tree decomposition of $G$ s.t. it is $k$-sparse with respect to $\mathcal{H}$, then there is an intersection support for $(G,\mathcal{H,K})$ of treewidth at most $k$.
Proof.
We construct a tree decomposition $(T,\mathcal{B}^{\prime})$ of a graph $\tilde{Q}$. For each bag $B\in\mathcal{B}$ and each subgraph $K\in\mathcal{K}$ intersecting $B$, we put a vertex in bag $B^{\prime}\in\mathcal{B}^{\prime}$ for each $H\in\mathcal{H}_{K}$ s.t. $H$ intersects $B$, where $B^{\prime}$ is the bag in $(T,\mathcal{B}^{\prime})$ corresponding to the bag $B$ in $(T,\mathcal{B})$ Since $(T,\mathcal{B})$ is $k$-sparse, there are at most $k$ vertices of $\mathcal{H}$ in $B^{\prime}$. We put a complete graph on the vertices of $\mathcal{H}$ in $B^{\prime}$ and obtain a tree decomposition $(T,\mathcal{B}^{\prime})$ of a graph $\tilde{Q}$ on $\mathcal{H}$. Since each $H\in\mathcal{H}$ is connected, the vertex in $\tilde{Q}$ corresponding to $H$, lies in a connected set of bags of $(T,\mathcal{B}^{\prime})$. Further, by construction, each edge between a pair of vertices $H,H^{\prime}$ of $\tilde{Q}$ lies in some bag of $\mathcal{B}^{\prime}$. Hence, $(T,\mathcal{B}^{\prime})$ is a valid tree decomposition of $\tilde{Q}$.
We now show that $\tilde{Q}$ is an intersection support. Consider any $K\in\mathcal{K}$. Since $K$ is a connected subgraph of $G$, the vertices of $K$ lie in a sub-tree of $T$. Since $(T,\mathcal{B})$ is $\mathcal{K}$-easy, for each adhesion set intersected by $K$, there is a subgraph $H\in\mathcal{H}_{K}$ intersecting that adhesion set. This implies that $\mathcal{H}_{K}$ induces a connected subgraph of $\tilde{Q}$, and hence $\tilde{Q}$ is an intersection support for $(G,\mathcal{H},\mathcal{K})$. ∎
We now show how we can modify any non-piercing system $(G,\mathcal{H,K})$ to one satisfying the conditions of Lemma 33.
See 9
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. Suppose $(T,\mathcal{B})$ is not $\mathcal{K}$-easy. We define $\phi:V(G)\to\{\mathbf{r},\mathbf{b}\}$. For each $v\in V(G)$, if $\mathcal{K}_{v}\neq\emptyset$ and $\mathcal{H}_{v}\neq\emptyset$, we set $\phi(v)=\mathbf{b}$. Otherwise, set $\phi(v)=\mathbf{r}$. Under this coloring, since $(T,\mathcal{B})$ is not $\mathcal{K}$-easy, there is a subgraph $K\in\mathcal{K}$ and an adhesion set $A$ s.t. $K\cap A\subseteq\mathbf{r}(A)$. Since the subgraphs in $\mathcal{K}$ are non-piercing, by Lemma 30, we obtain a tree decomposition $(T,\mathcal{B}^{\prime})$ of treewidth $t^{\prime}=2^{t+2}+t$ that is easy with respect to $\mathcal{K}$. By the choice of coloring, it implies $(T,\mathcal{B}^{\prime})$ is $\mathcal{K}$-easy.
Each $K\in\mathcal{K}$ induces a connected subgraph of $G$. Hence, $K$ intersects a connected set of bags of $(T,\mathcal{B^{\prime}})$. Since $(T,\mathcal{B}^{\prime})$ is $\mathcal{K}$-easy, for any pair of vertices $u,v\in K$, there is a path in $T$ between a bag containing $u$ and a bag containing $v$ s.t. each adhesion set on this path is intersected by a subgraph $H\in\mathcal{H}_{K}$.
Recall that to obtain dual support for $(G,\mathcal{H})$, the Algorithm $k$-SDS adds a complete graph on the subgraphs in $\mathcal{H}$ intersecting each bag of $(T,\mathcal{B}^{\prime})$. Therefore, a dual support for $(G,\mathcal{H})$ thus obtained, is also an intersection support for $(G,\mathcal{H},\mathcal{K})$. However, this support may not have bounded treewidth. To obtain a support of small treewidth, we need to first sparisfy the subgraphs in $\mathcal{H}$.
Since the subgraphs in $\mathcal{H}$ are non-piercing, by Lemma 32, we can obtain a collection $\mathcal{H}^{\prime}$ s.t. $(T,\mathcal{B}^{\prime})$ is $2^{4(t^{\prime}+1)}$-sparse with respect to $\textrm{unique}(\mathcal{H}^{\prime})$. Note that $(T,\mathcal{B}^{\prime})$ remains $\mathcal{K}$-easy w.r.t. the intersection system $(G,\textrm{unique}(\mathcal{H}^{\prime}),\mathcal{K})$. Now, by Algorithm $k$-SDS we obtain a tree decomposition of a graph $Q^{*}$ that by Lemma 31 is a dual support for $(G,\textrm{unique}(\mathcal{H}^{\prime}))$ and $\textsc{tw}(Q^{*})\leq 2^{4(t^{\prime}+1)}=2^{{2^{t+4}}+4(t+1)}$. Since $Q^{*}$ is obtained by $k$-SDS, it also an intersection support for $(G,\textrm{unique}(\mathcal{H}^{\prime}),\mathcal{K})$. By Theorem 8, we can extend $Q^{*}$ to a dual support $\tilde{Q}$ for $(G,\mathcal{H})$ such that $\textsc{tw}(\tilde{Q})=\textsc{tw}(Q^{*})$. Then $\tilde{Q}$ is also an intersection support for $(G,\mathcal{H},\mathcal{K})$ since the tree decomposition $(T,\mathcal{B}^{\prime})$ is $\mathcal{K}$-easy.
6.4 Implementation
In this section, we show that our algorithms for the construction of supports of bounded treewidth run in polynomial time if the treewidth of the host graph is bounded. We argue by showing a result below that generalizes Theorem 22 to graphs of treewidth $t$. We show that if the treewidth of $G$ is bounded, then $|\mathcal{H}|$ and $|\mathcal{K}|$ are bounded by $\textrm{poly}(n)$ where $n=|V(G)|$.
Theorem 34.
Let $(G,\mathcal{H},\mathcal{K})$ be a non-piercing intersection system of treewidth $t$. Then $|\mathcal{H}|,|\mathcal{K}|=O(n^{3t+3})$.
Proof.
We show that $|\mathcal{H}|=O(n^{3t+3})$. The bound on the size of $\mathcal{K}$ follows analogously.
First, we will show that the VC-dimension of the set system defined by $(V(G),\mathcal{H})$ is at most $33t+$ 3. Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. We will show that no set of size more than $33t+$ 3 can be shattered. Suppose not. Let $S$ be a set with $43t+$ 4 vertices that can be shattered. There is an edge $\{x,y\}$ in $T$ such that the union of bags in $T_{x}$ contains a subset $S_{1}\subset S$ of at least $t+2$ vertices, and the union of bags in $T\setminus T_{x}$ contains a subset $S_{2}\subset S$ of at least $t+2$ vertices where $S_{1}\cap S_{2}=\emptyset$. Note that such subsets should exist since each bag contains at most $t+1$ vertices and $|S|=3t+4$.
There are at least $k=\binom{t+2}{\lfloor\frac{t+2}{2}\rfloor}$ subsets of $S_{i}$ of size $\lfloor\frac{t+2}{2}\rfloor$ for $i\in\{1,2\}$. Let $H_{1},H_{2},\ldots H_{k}$ be the distinct subgraphs in $\mathcal{H}$ defined by these $k$ sets in $S_{1}$, and $H^{\prime}_{1},H^{\prime}_{2},\ldots H^{\prime}_{k}$ be the distinct subgraphs in $\mathcal{H}$ defined by these $k$ sets in $S_{2}$. Note that $H_{i}$’s are pairwise incomparable in the containment order. Similarly, $H^{\prime}_{i}$ ’s are pairwise incomparable.
Next, we define a set of subgraphs in $\mathcal{H}$ that intersect with both $S_{1}$ and $S_{2}$. Such a set of subgraphs must exist since $S$ is shattered by element of $\mathcal{H}$. Let $J_{i}\in\mathcal{H}$ be a subgraph such that $J_{i}\cap S=H_{i}\cup H^{\prime}_{i}$ for $i=1,2,\ldots k$. Then, each $J_{i}$ contains exactly $2\lfloor\frac{t+2}{2}\rfloor$ vertices of $S$ and hence, they are pairwise incomparable in $S$. Also, each $J_{i}$ intersects the adhesion set $A_{e}$ as they are connected subgraphs. Since the subgraphs in $\mathcal{H}$ are non-piercing, by part $(ii)$ of Lemma 27, $J_{i}$ and $J_{\ell}$ should be incomparable in $A_{e}$ for $i\neq\ell$. But there are only $\binom{t+1}{\lfloor\frac{t+1}{2}\rfloor}<k$ subsets of $A_{e}$ that are pairwise incomparable. This contradicts the fact that the set $S$ can be shattered. Hence, the VC-dimension of $(V(G),\mathcal{H})$ is at most $33t+$ 3. By Sauer-Shelah lemma (see [Mat02LecDiscGeom], Chap. 10) therefore, $|\mathcal{H}|\leq O(n^{3t+3})$.
Now, we are ready to show the running time of our algorithms.
Theorem 35.
Let $G$ be an $n$-vertex graph of treewidth $t$ and $\mathcal{H},\mathcal{K}$ be non-piercing subgraphs of $G$. Then, a primal and a dual support of treewidth $O(2^{t})$ and $O(2^{4t})$ respectively can be computed in time $\textrm{poly}(n^{t})$. And, an intersection support of treewidth at most $2^{O(2^{t})}$ can be computed in time $\textrm{poly}(n^{2^{t}})$.
Proof.
By Theorem 7, 8 and 9, there are appropriate supports of the claimed treewidth. We show below the running time of the algorithms.
Since $\textsc{tw}(G)=t$, a tree decomposition $(T,\mathcal{B})$ of width $t$ can be computed in time $2^{O(t)}n$ where $T$ is a binary tree and has $O(n)$ number of nodes [bodlaender1998partial, bodlaender2016, BODLAENDER1996358].
For the construction of a primal support, we require an easy tree decomposition of $G$. We did a post-order traversal of the edges in $T$ and chose a minimal subgraph in $\mathcal{M}_{S}$ for each non-empty subset $S$ of an adhesion set $A$ in $T$. Since $A$ has at most $t+1$ vertices, there are at most $2^{t+1}$ choices for $S$. Also, by Theorem 34, $|\mathcal{H}|=n^{3t+3}$. Therefore, a primal support of treewidth $2^{t+2}+t$ can be computed in time $\textrm{poly}(n^{t})$.
For a dual support, we first construct a graph system that is $2^{4(t+1)}$-sparse. For such a construction, we again do a post-order traversal of the edges in $T$ and choose a subset $S$ of an adhesion set $A$ to select a pusher in $\mathcal{M}_{S}$. In $O(|\mathcal{H}|^{2})$ time, we can do the pushing operation at a subset $S$ to get new subgraphs. Therefore all the adhesion sets can be processed in time $O(2^{t}n|\mathcal{H}|^{2})$. Once a $2^{4(t+1)}$-sparse tree decomposition is obtained, the algorithm $k$-SDS puts a complete graph on the subgraphs in each bag and hence can be done in $O(k^{2}t)$ time where $k=2^{4(t+1)}$. Therefore, a dual support of treewidth $2^{4(t+1)}$ can also be computed in time $\textrm{poly}(n^{t})$ since $|\mathcal{H}|\leq n^{3t+3}$ by Theorem 34.
Finally, the construction of an intersection support follows the construction of primal and dual supports. By the arguments above, it follows that a $\mathcal{K}$-easy tree decomposition can be computed in time $\textrm{poly}(n^{t})$ where the width of the resulting tree decomposition is $O(2^{t})$. Then a $2^{O(2^{t})}$-sparse tree decomposition can be computed in time $\textrm{poly}(n^{2^{t}})$. Hence, the total running time to compute an intersection support is $\textrm{poly}(n^{2^{t}})$. ∎
7 Lower Bounds
In this section, we show that there exist graphs of treewidth $t$ whose (primal or dual) support requires treewidth $\Omega(2^{t})$.
Theorem 36.
There is a non-piercing graph system $(G,\mathcal{H})$ with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ s.t. $\textsc{tw}(G)\leq n+1$, but $\textsc{tw}(Q)\geq 2^{\lceil n/2\rceil}$ for any primal support $Q$.
Proof.
Let $m=\lceil n/2\rceil$ and $N=2^{m}$. We construct a graph $G$ and non-piercing subgraphs $\mathcal{H}$ as follows. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of isolated vertices colored $\mathbf{b}$ such that there is no edge between the vertices of $B$. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices colored $\mathbf{r}$. Let $\textrm{bin}(x)$ denote the binary representation of $x\in\mathbb{N}$. By $\textrm{bin}(x)$ of an indexed set $X$, we mean the subset of $X$ with indices corresponding to $1$s in $\textrm{bin}(x)$. Since $\log N=m$, for any integer $0\leq x\leq N$, $\textrm{bin}(x)$ of $C$ and $\textrm{bin}(x)$ of $R$ are well-defined.
For a horizontal pair of grid vertices $e=\{b_{i,j},b_{i,j+1}\}$, we connect these vertices to $\textrm{bin}(j)$ of $C$, $\textrm{bin}(j+1)$ of $C$ and $\textrm{bin}(i)$ of $R$. This gives an induced subgraph $H_{i,j,i,j+1}$. Similarly, for a vertical pair $e^{\prime}=\{b_{i^{\prime},j^{\prime}},b_{i^{\prime}+1,j^{\prime}}\}$, we connect these vertices to $\textrm{bin}(i^{\prime})$ of $R$, $\textrm{bin}(i^{\prime}+1)$ of $R$, and $\textrm{bin}(j^{\prime})$ of $C$. This gives us an induced subgraph $H_{i^{\prime},j^{\prime},i^{\prime}+1,j^{\prime}}$. This completes simultaneously, the construction of $G$ and the subgraphs $\mathcal{H}$.
We claim that $(G,\mathcal{H})$ is a non-piercing graph system of treewidth at most $n+1$. The graph $G$ thus constructed is a bipartite graph with bipartition $(B,R\cup C)$, since the vertices in $B$ are pairwise non-adjacent, and the vertices in $R\cup C$ are pairwise non-adjacent. Since $|R\cup C|=2\left\lceil n/2\right\rceil$, it implies that $\textsc{tw}(G)\leq n+1$.
The subgraphs in $\mathcal{H}$ are connected. Let $H,H^{\prime}\in\mathcal{H}$ be arbitrary. Since they correspond to distinct pairs of points of the grid, they differ in at least one grid vertex, and they also differ in their adjacency in either $R$ or in $C$. Hence, $H\setminus H^{\prime}$ is connected via the grid vertex in $H\setminus H^{\prime}$. Similarly, $H^{\prime}\setminus H$ is connected. Therefore, $\mathcal{H}$ is a collection of non-piercing subgraphs of $G$.
Each $H\in\mathcal{H}$ contains exactly two vertices of $B$, which should be adjacent in any primal support for $(G,\mathcal{H})$. Therefore, any primal support $Q$ must contain a grid graph on $B$. Hence, $\textsc{tw}(Q)\geq N$.
Theorem 37.
There is a non-piercing graph system $(G,\mathcal{H})$ s.t. $\textsc{tw}(G)\leq n+1$ and for any dual support $Q^{*}$, $\textsc{tw}(Q^{*})\geq 2^{\lceil n/2\rceil}-1$.
Proof.
The construction is similar to the construction of the primal system. Let $m=\lceil n/2\rceil$ and $N=2^{m}$. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of isolated vertices. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices, none of which are pairwise adjacent.
The graph system $(G,\mathcal{H})$ is constructed as follows. A face $F_{i,j,i+1,j+1}$ for $i,j\in\{1,\ldots,N-1\}$ is the subset $\{b_{i,j},b_{i,j+1},b_{i+1,j},b_{i+1,j+1}\}$ of $B$. For each $F_{i,j,i+1,j+1}$, we connect all its vertices to $\textrm{bin}(i)$ and $\textrm{bin}(i+1)$ of $R$, and $\textrm{bin}(j)$ and $\textrm{bin}(j+1)$ of $C$. Finally, for every two consecutive vertices on the grid, we put a vertex in the middle. Let $B^{\prime}$ be the set of vertices added. Thus, $B\cup B^{\prime}$ forms a grid of size $(2N-1)\times(2N-1)$. For each vertex of $B^{\prime}$ that lies between two consecutive vertices $x,y$ of $B$ row-wise or column-wise, we make it adjacent to all the vertices in $R\cup C$ that are adjacent to both $x$ and $y$. This completes the construction of $G$. Now, for each face $F_{i,j,i+1,j+1}$, we add an induced subgraph $H_{i,j,i+1,j+1}$ to $\mathcal{H}$ consisting of the vertices of $F_{i,j,i+1,j+1}$, the four vertices of $B^{\prime}$ that lie between the vertices of this face, and the common neighbours of these 8 vertices in $R\cup C$ i.e., $\textrm{bin}(i)$ and $\textrm{bin}(i+1)$ of $R$, and $\textrm{bin}(j)$ and $\textrm{bin}(j+1)$ of $C$. This results in connected induced subgraphs $\mathcal{H}$ of $G$.
$G$ is a bipartite graph with bipartition $(B,C\cup R)$. Since $|C\cup R|=2m$, $\textsc{tw}(G)\leq 2m$. Each subgraph in $\mathcal{H}$ contains 8 vertices of $B\cup B^{\prime}$. Since any two subgraphs $H,H^{\prime}\in\mathcal{H}$ share at most 3 vertex of $B\cup B^{\prime}$ along a row or column of the grid. Hence, $H\setminus H^{\prime}$ is connected via the vertices of $B\cup B^{\prime}$ in $H\setminus H^{\prime}$. Thus, $\mathcal{H}$ is a collection of non-piercing subgraphs.
Since each vertex in $B^{\prime}$ lies in exactly two subgraphs, the subgraphs in $\mathcal{H}$ form a grid of size $(N-1)\times(N-1)$. Therefore, for any dual support $Q^{*}$, $\textsc{tw}(Q^{*})\geq N-1$.
8 Conclusion
In this paper, we gave algorithms to construct support for hypergraphs defined on a host graph $G$ and collections of non-piercing subgraphs of $G$. When $G$ is outerplanar, we showed that there are outerplanar primal, dual and intersection supports. Deciding if an abstract hypergraph admits a 2-outerplanar support was shown to be NP-hard by Buchin et al. [buchin2011planar]. The complexity of the decision problem is not known for outerplanar support. It is plausible that the machinery we used for the construction of outerplanar support may help in resolving this open problem.
For the case when $G$ has treewidth $t$, we constructed primal, dual and intersection supports of treewidth at $O(2^{t})$, $O(2^{4t})$ and $2^{O(2^{t})}$, respectively. All our algorithms run in polynomial time in the number of vertices of $G$ if $t$ is bounded above by a constant. We also construct examples where the exponential blow-up in the treewidth of any primal or dual supports can not be improved. In the case of dual setting, we constructed a support of treewidth $O(2^{4t})$. We leave it an open question if the exponent $4t$ can be improved to $t$. Similarly, in the intersection setting, there is a double exponential in the treewidth of an intersection support. We believe that it is possible to get an intersection support of treewidth $O(2^{ct})$ for some constant $c$.
From our results, it follows that a non-piercing outerplanar graph system admits an outerplanar support. Since $\textsc{tw}(G)=2$ for an outerplanar graph $G$, we wonder if for a non-piercing graph system of treewidth 2, there exists a primal/dual or intersection support of treewidth 2.